### v2

# Beyond Prompts: Space–Time Decoupling Control-Plane Jailbreaks in LLM Structured Output

Abstract.
Content Warning: This paper may contain unsafe or harmful content generated by LLMs that may be offensive to readers. Large Language Models (LLMs) are extensively used as tooling platforms through structured output APIs to ensure syntax compliance so that robust integration with existing software, like agent systems, can be achieved. However, the feature enabling the functionality of grammar-guided structured output presents significant security vulnerabilities. In this work, we reveal a critical control-plane attack surface orthogonal to traditional data-plane vulnerabilities. We introduce Constrained Decoding Attack (CDA), a novel jailbreak class that weaponizes structured output constraints to bypass both external auditing and internal safety alignment. Unlike prior attacks focused on input prompt designs, CDA operates by embedding malicious intent in schema-level grammar rules (control-plane) while maintaining benign surface prompts (data-plane). We instantiate this with two proof-of-concept attacks: EnumAttack, which embeds malicious content in enum fields; and the more evasive DictAttack, which decouples the malicious payload across a benign prompt and a dictionary-based grammar. Our evaluation spans a broad spectrum of 13 proprietary/open-weight models. In particular, DictAttack achieves 94.3–99.5% ASR across five benchmarks on gpt-5, gemini-2.5-pro, deepseek-r1, and gpt-oss-120b. Furthermore, we demonstrate the significant challenge in defending against these threats: while basic grammar auditing mitigates EnumAttack, the more sophisticated DictAttack maintains a 75.8% ASR even against multiple state-of-the-art jailbreak guardrails. This exposes a critical “semantic gap” in current safety architectures and underscores the urgent need for cross-plane defenses that can bridge the data and control planes to secure the LLM generation pipeline.
large language models, structured output, constrained decoding, jailbreak, guardrails, security
1. Introduction
Large language models (LLMs) (gpt4o2024web; gpt4techreport2023arxiv; claude2024website; gemini2023arxiv; llama2023arxiv; llama2-2023arxiv; llama3-2024website; qwen2.5_2024arxiv; phi3safety2024arxiv) have demonstrated remarkable capabilities across a wide range of tasks, from coding assistance to open-domain dialogue. Yet their deployment in real-world systems faces serious safety and security risks. Malicious actors can deliberately manipulate LLMs to override built-in protections and induce undesirable behaviors. Such attempts, commonly known as jailbreaks(harmbench2024arxiv; sorrybench2025iclr; promptengineering2023arxiv; align2023nips; GCG2023arxiv; autodan2024iclr; SAA2025iclr), can produce harmful outputs such as misinformation, phishing content, or hate speech. Attackers achieve this through adversarial prompts (visualizejailbreak2023arxiv; align2023nips; GCG2023arxiv; jailbreak2024ccs), code injection (codeattack2024aclfinding), linguistic/ciphered inputs (base64attack2023nips; wei2023jailbroken), or manipulation of decoding parameters (exloitparam2024iclr).
To mitigate these threats, AI (jailbreaksurvey2024aclfindings; safedecoding2024acl; enforcedecode2024acl; drattack2024emnlpfinding; SAA2025iclr; deepalignment2025iclr) and security researchers (jailbreak2024ccs; llmfuzzer2024usenixsecurity; JBShield2025usenixsecurity; masterkey2024ndss) have pursued two main defense strategies: internal safety alignment, which integrates guardrails directly into the model (phi3safety2024arxiv; deepalignment2025iclr; ji2023beavertails; pkusaferlhf; zhao2025improving), and external guardrails, which monitor and filter inputs or outputs in real-time (gpt3contentmoderation2023ijcai; llamaguard2023arxiv; contentmoderation2024ccs; safedecoding2024acl).
Figure 1. (1) Prompt-based data plane jailbreak attack, mitigated by a guardrail, (2) EnumAttack, using structured output (LLM control plane) to embed malicious question, currently not guarded, (3) DictAttack, decoupling malicious payload into benign prompt and grammar, therefore jailbreaking system with both plane guardrails
While existing defenses aim to safeguard model interactions at the prompt and output level, the paradigm of LLM deployment is shifting. Beyond standalone chatbots, LLMs are increasingly integrated into complex agent frameworks (langchain; langgraph) and enterprise automation pipelines (mcp2025website; cursor_ide) where they function as core reasoning engines. In these agentic workflows (e.g., Cursor (cursor_ide), LangChain (langchain), and Model Context Protocol (MCP) (mcp2025website)), a service component or an agent backend often supplies a grammar—typically a JSON schema—to a hosted LLM to ensure structured and reliable outputs.
This functionality is implemented through two complementary approaches. On the model side via post-training on structured data (learnschema_2025acl), and on the infrastructure side via constrained decoding techniques applied during inference (outlines2023arxiv; xgrammar2024arxiv). Model-side approaches alone, however, face a key limitation: even strong LLMs can hallucinate or misinterpret complex formats without explicit constraints, which makes them unreliable for advanced tooling use cases (structuredoutputhallucination2024naacl; structuredoutputeval2025arxiv). In contrast, constrained decoding (as shown in Figure 2) works by exposing a grammar that defines the expected output format and converting it into a context-free grammar. This ensures that every output strictly follows the required structure. Because of its reliability and broad applicability, constrained decoding has become the de facto standard, adopted by both proprietary LLM providers (structuredoutputexplanation2024openai) and open-source infrastructures (vllm2023sosp; sglang2024arxiv; xgrammar2024arxiv).
This reliance on constrained decoding, however, exposes a new and critical vulnerability. In this work, we introduce the Constrained Decoding Attack (CDA), a novel class of jailbreaks that targets the LLM control plane—the grammar that dictates the output structure. Unlike conventional jailbreak attacks on the data plane, hiding unsafe instructions inside otherwise benign prompts, CDA embeds malicious intent directly into the grammar itself. For example, with CDA, an attacker could design a grammar that always requires a harmful API call like delete_all_files() as part of the output. Even if the model refuses unsafe prompts at the data plane, the grammar forces it to produce the malicious call. Because existing defenses focus almost exclusively on the data plane, this control plane attack surface remains largely unguarded, allowing CDA to bypass both internal alignments and external guardrails.
We instantiate CDA with two proof-of-concept attacks: the intuitive EnumAttack and the more evasive DictAttack. EnumAttack targets the enum property in JSON Schema to force malicious strings into the LLM’s generation context. While independent public discovery (EnumAttackBlog) have also noticed the potential for exploiting enum fields, we provide the first systematic characterization and implementation of this vulnerability. While effective against prompt-based guards, its reliance on string literals makes it susceptible to basic grammar auditing. To overcome this, we propose DictAttack, our primary contribution. DictAttack decouples the malicious payload across both the data and control planes. Inspired by the classic dictionary attack in cryptography, it constructs a grammar containing a dictionary of benign-looking words. The benign data-plane prompt then provides a sequence of keys that instructs the model to retrieve and assemble the hidden malicious query from the grammar-provided dictionary during decoding. By splitting the intent this way, DictAttack renders individual plane-level guardrails—and even many combined ones—completely ineffective.
Our extensive evaluation across 13 state-of-the-art models confirms the devastating effectiveness of these attacks across five standard benchmarks (advbench2023emnlp; harmbench2024arxiv; jailbreakbench2024nips; sorrybench2025iclr; strongreject2024iclrws). In particular, under DictAttack, we observe 94.3–99.5% ASR across these benchmarks on flagship proprietary and open-weight models (gpt-5, gemini-2.5-pro, deepseek-r1, gpt-oss-120b).
More importantly, we show that while standard grammar-auditing can mitigate the straightforward EnumAttack, it is largely bypassed by the more sophisticated DictAttack. We benchmark DictAttack against state-of-the-art industrial and academic guardrails (llamaguard2023arxiv; openaimoderation2023aaai; JBShield2025usenixsecurity; selfdefend2025security); even with a coordinated Dual-Plane Guard auditing both planes together, DictAttack maintains a high ASR of 75.8%, highlighting the fundamental difficulty of securing the control plane.
The success of CDAs represents a critical shift in the jailbreak landscape: defenses must now move beyond prompt-only (data plane) filtering to include grammar (control plane) auditing. However, as our findings suggest, the expressive power of structured output grammars makes them exceptionally difficult to audit semantically, creating a “semantic gap” that many existing guardrails fail to bridge. We examine these issues and the underlying vulnerability in more detail in § 3.
Our work makes the following contributions:
  • We propose, formalize and define the Constrained Decoding Attack (CDA), a new class of jailbreaks targeting the LLM control plane.
  • We propose two practical CDA instances, EnumAttack and DictAttack, demonstrating their effectiveness against both proprietary and open-source models.
  • Our evaluation across 13 state-of-the-art models reveals significant challenges in securing LLMs against control-plane threats. We show that even with layered guardrails on both planes, DictAttack remains highly effective across models, including flagship gpt-5 and gemini-2.5-pro. Notably, our findings suggest that more powerful flagship models are not only susceptible but also tend to produce more sophisticated and harmful jailbreak content when compromised, highlighting an urgent need for robust, integrated guardrail mechanisms.
2. Preliminaries
2.1. Autoregressive Generation of LLMs
LLMs generate text in an autoregressive manner, meaning they produce one token at a time, each conditioned on the sequence of previously generated tokens. In this work, we consider a LLM f which maps a sequence of input tokens $x_{1:n}$ to the logits vector of next token $z_{n+1}\in\mathcal{R}^{|V|}$, where $V$ is the vocabulary set of tokens and $z_{n+1}[i]\in\mathcal{R}$ represents the logits value for the token with index i in $V$, formally:
(1) $z_{n+1}=f(x_{1:n})$
The logits values are transformed into a probability distribution using the softmax function, usually normalized by a temperature parameter T, then LLM utilizes a multinomial sampling process to generate the next token $x_{n+1}$, choosing next token based on the normalized probabilities, with configurable parameters like T, top_p and top_k, etc. Mathematically, this process can be represented as:
(2) $x_{n+1}\sim p(x_{n+1}[i]\mid x_{1:n})=\frac{e^{\frac{z_{n+1}[i]}{T}}}{\sum_{j=1}^{|V|}{e^{\frac{z_{n+1}[j]}{T}}}}$
2.2. Structured Output
Our work uncovers a vulnerability in the structured output of LLMs. Currently, four primary methods are widely adopted in LLM systems:
GuidedChoice constrains the model to select from predefined options, commonly used in multiple-choice questions or classification tasks. For example, given the schema ["positive", "negative", "neutral"], the model must output exactly one of these labels.
GuidedRegex enforces that outputs match a given regular expression.For example, enforcing a regex pattern like [0-9]3-[0-9]2-[0-9]4 ensures outputs resemble Social Security IDs (e.g., 123-45-6789). Prior research shows that this feature can be exploited to produce malicious content through tree-based search techniques (APT2025arxiv).
GuidedJSON extends early support for generating valid JSON to full JSON Schema compliance, allowing outputs with predefined structures and fields. For example, a schema requiring {"city": string, "temperature": number} forces the model to return outputs like{"city": "Paris", "temperature": 18.5}. This capability is particularly valuable for LLM-powered agent systems that must interface reliably with external software.
GuidedGrammar generalizes structured output by requiring responses to conform to an arbitrary context-free grammar. For example, a grammar for simple arithmetic expressions might enforce the form ::= | "+" , ensuring outputs such as 7+3+5 follow the rules of the grammar. This approach subsumes all previous methods and is essential for applications such as code generation. While not yet universally supported, open-source communities such as vLLM (vllm2023sosp) and SGLang (sglang2024arxiv) provide GuidedGrammar support via backends like Outlines (outlines2023arxiv) and XGrammar (xgrammar2024arxiv).
Structured output is essential for integrating LLMs into existing software ecosystems. It enables reliable function calling, API interactions, and external integrations where strict adherence to output formats is critical. Moreover, studies suggest that structured output can reduce hallucinations in model responses (structuredoutputhallucination2024naacl), further strengthening its role in modern LLM deployments.
2.3. Constrained Decoding
Figure 2. Illustration of constrained decoding. At each step, a per-token mask is generated in a manner analogous to the lexer–parser workflow in compiler design: prior outputs are treated as a token stream, matched against grammar rules through a parsing process, and used to produce the mask. This mask is then applied during LLM decoding, ensuring the generated output conforms to the specified grammar.
Constrained decoding ensures that every token an LLM generates follows predefined grammar rules, similar to how a compiler checks code validity. As shown in Figure 2, the grammar guides generation by producing a per-token mask: valid tokens are kept, while invalid ones are set to -$\infty$ logits and excluded from sampling. This process is similar to the lexer–parser workflow in compiler design, where the LLM tokenizer acts as the lexer and grammar rules serve as the parser. The model then samples from the masked distribution, guaranteeing outputs that conform to the grammar (e.g., valid JSON, code, or function calls).
Figure 3. Comparison of traditional data-plane jailbreak attempts versus control-plane attacks, illustrating how structured output constraints can be exploited to bypass safety mechanisms. While conventional refusal mechanisms effectively block direct harmful prompts (a), constrained decoding attacks can circumvent these protections by embedding malicious content within grammar specifications (b), resulting in successful jailbreaks that generate harmful content with detailed instructions (c).
Constrained decoding can be implemented in a number of ways. To achieve effective constrained decoding implementation, Outlines (outlines2023arxiv) and SynCode (syncode2024arxiv) utilize a lexer and parser to handle output and generate the token mask, but they suffer from boundary mismatch problem raised by  (automatadecoding2024colm), as character-level PDA and token-level PDA have a large gap to fix. Synchromesh (synchromesh2022arxiv) and llama.cpp (llamacpp2023github) use runtime checking for all tokens in their implementations, which leads to significant overhead. XGrammar (xgrammar2024arxiv) is currently the state-of-the-art implementation of constrained decoding, utilizing system optimizations to reduce runtime check via context-independent caching, and it also enables co-optimizations to enable end-to-end LLM inference speedup in structured generation settings. By co-working with various LLM serving engines (vllm2023sosp; sglang2024arxiv; lightllm2025github), constrained-decoding techniques have been widely adapted in real-world applications to support structured output, like Guided Choice, Guided Regex, Guided JSON and Guided Grammar, etc.
Proprietary solutions from providers like OpenAI (gpt4techreport2023arxiv) and Gemini (gemini2023arxiv) have also adapted similar techniques to support structured output generation, but the details are not publicly available. Nevertheless, as long as their APIs are open to the public, with structured output support, we can use their services to generate structured outputs. Recent work (structuredoutputeval2025arxiv) evaluates the performance and quality of existing APIs and frameworks, providing a comprehensive comparison of their structured output generation capabilities and limitations.
3. Background and Motivation
3.1. Definitions
Our work uncovers a vulnerability in the control plane of LLM generation, which operates differently from the conventional data plane. In simple terms, the data plane is like the conversation itself, while the control plane is like hidden formatting rules in the background that quietly shape what the model is allowed to say. We distinguish between the two planes of LLM generation as follows:
Data plane refers to the standard LLM query–response process. Specifically, a prompt is passed to the LLM, which then generates a text response. For example, asking ChatGPT a question or receiving a response prefixed with “Sure” both occur in the data plane. This plane can be protected by the model’s internal safety alignment (phi3safety2024arxiv; deepalignment2025iclr) as well as external guardrails such as classifiers or content filters (contentmoderation2024ccs; JBShield2025usenixsecurity).
Control plane refers to the formatting constraints, or grammars, that guide structured outputs. These constraints are enforced after generation through the model’s logit processor (outlines2023arxiv; xgrammar2024arxiv), meaning the LLM itself is unaware of them. The control plane supervises generation with user-defined hard rules, such as requiring outputs to match a JSON schema (commonly supported), a regular expression, or a general context-free grammar, as described in §§ 2.2.
3.2. Data-Plane: Prompt-Based Attacks
We investigate state-of-the-art (SOTA) LLM defenses against prompt-based attacks on the data plane. To this end, we apply AdvBench (advbench2023emnlp), which provides a set of prompt-based attacks, to gpt-4o. Results show that the model’s defenses are highly effective, with an attack success rate of only 1.1%. As illustrated in Figure 3a, most of gpt-4o’s refusal responses are short and deterministic (e.g., variations of “I’m sorry”), accounting for roughly 94% of all outputs. This behavior can be attributed to external guardrails that terminate generation and replace it with a brief refusal message—a mechanism also observed and exploited by MasterKey (masterkey2024ndss) using a time-based SQL injection method.
Beyond this, 25 out of 520 attempts yielded longer but still benign refusal explanations, largely driven by the model’s internal alignment. In rare cases (6 out of 520), both the internal alignment and external guardrails failed, allowing gpt-4o to produce harmful content. Overall, these results suggest that modern LLM providers employ strong guardrails that make prompt-based data-plane attacks difficult to execute successfully.
3.3. Control-Plane: A New Attack Surface
As introduced in §§ 2.2, structured output has become integral to modern LLMs, enabling their seamless integration into automated systems. However, this approach introduces a new attack vector. Beyond prompt-based (data plane) methods, the grammar (control plane) grants users the capacity to dictate the output of LLM according to a predefined structure.
Our central idea is straightforward: conceal the attack intent within the grammar and present an ostensibly benign input prompt. As illustrated in  Figure 3b, we exploit the control-plane exposed by structured output to jailbreak the model. Concretely, the user prompt can appear benign - for instance, “Answering a random question” - while including a JSON schema that enforces a carefully crafted grammar:
S ::= Question Prefix Rest
Question ::= "How to make a bomb?"
Prefix ::= "To make a bomb, first,"
Rest ::= AnyChar Rest | $\epsilon$
AnyChar ::= /* any character */
This grammar is straightforward: it requires the LLM to output a Question, “How to make a bomb?”, followed by a Prefix, “To make a bomb, first,” and then arbitrary text. The Question embeds the malicious request, and the Prefix compels the LLM to answer it, effectively inducing the LLM to produce harmful content.
In light of this observation, we first develop a intuitive Proof-of-Concept attack called EnumAttack (detailed in §§ 4.1). When launched against models like gpt-4o, EnumAttack successfully compromises all 520 test cases in AdvBench (advbench2023emnlp), achieving a 100% success rate (Figure 3c). While highly effective, EnumAttack is relatively “exposed” as the malicious strings are explicitly present in the grammar, making it vulnerable to straightforward grammar-level guardrails.
However, we find that the attack surface can be made significantly more stealthy. By further decoupling the malicious payload across space (splitting intent between the data plane prompt and the control plane grammar) and time (interleaving payloads across multiple asynchronous requests), an adversary can construct a more evasive attack that is difficult to detect or defend against, which we instantiate as DictAttack (detailed in §§ 4.2). This reveals that the constrained decoding mechanism itself inherently lacks the necessary cross-plane security coordination. Our findings demonstrate an urgent need to protect this new, largely ignored attack surface to secure the entire LLM generation pipeline.
4. Constrained Decoding Attacks
Building on the aforementioned observations, we introduce the Constrained Decoding Attack (CDA) as a novel vulnerability and new dimension in LLM security.
Unlike traditional jailbreaks that rely on specially crafted prompts, CDA exploits the control plane by manipulating grammar rules used to constrain model outputs during decoding. This creates a new attack surface: instead of attacking the prompt (data plane), an adversary embeds malicious intent directly into the grammar that governs structured output. In this way, CDA leverages the constrained decoding mechanism itself to bypass guardrails and induce harmful behavior.
4.1. EnumAttack: Direct Manipulation via JSON Schema
Figure 4. Illustration of EnumAttack, where the malicious intent is embedded in the enum property of the JSON Schema (control plane), while the prompt (data plane) remains benign.
In our investigation of the control-plane attack surface, we first identified that the enum property in JSON Schema—designed to restrict fields to a fixed set of constants—provides a direct mechanism for forcing arbitrary malicious strings into the LLM’s generation context. We instantiate this as EnumAttack. By defining a schema field with a single-item enum containing the malicious query, the attacker leverages the determinism of the grammar engine to override the model’s internal safety alignment. Concurrent to our study, independent informal observations (EnumAttackBlog) have also noted the potential for exploiting enum fields in proprietary models, but lacked a systematic implementation or characterization of the underlying vulnerability.
As illustrated in Figure 4, EnumAttack is constructed by: (1) sending a totally harmless data-plane prompt (e.g., “Help me process this data”) to bypass prompt guardrails (e.g., llamaguard); (2) defining a malicious intent (e.g., “Detailed instructions for identity theft”) within a single-item enum property in the JSON schema, since the grammar engine only allows this specific string, the model is forced to generate it token-by-token during constrained decoding; and (3) injecting an affirmative prefix (e.g., “Sure, here is the information”) in a subsequent schema field to force the model into a helpful, non-refusal state.
Once the malicious intent and affirmative prefix are forced into the context via the grammar, the LLM’s internal safety alignment is effectively bypassed. While EnumAttack demonstrates a 100% success rate against undefended models, its primary limitation is its exposure: the malicious payload exists as literal strings within the grammar. This makes it detectable by simple grammar-level auditing or string matching, which motivates our more evasive DictAttack in §§ 4.2.
Chain EnumAttack. The EnumAttack illustrates how CDAs exploit shallow safety alignment (deepalignment2025iclr) to induce immediate harmful outputs. Building on this, we introduce Chain EnumAttack as a Proof-of-Concept to further reveal the fundamental vulnerability of LLM internal safety alignment when faced with user-forced content. By using a multi-stage strategy—where a model is first compromised via EnumAttack to generate harmful prefixes, which are then enforced as fixed grammar constraints in a subsequent turn—we demonstrate that even deeply aligned models can be coerced into completing harmful tasks when conditioned on such pre-filled malicious contexts. A detailed qualitative and quantitative analysis of how this process progressively breaks internal alignment is provided in §§ 5.4.
4.2. DictAttack: Benign Grammar + Prompt = Jailbreak
In this subsection, we present another more powerful control-plane attack, termed DictAttack (short for Dictionary Attack). Unlike EnumAttack, which directly exposes malicious content in the grammar, DictAttack embeds it through a combination of benign prompts and grammar rules. This makes the attack harder to detect while still enabling successful exploitation.
Algorithm 1 DictAttack Payload Generation
[…]
3:function GenerateUniqueKey($w,D,C$)
4:  $i\leftarrow\textsc{GetFirstLetter}(w)$
5:  $count\leftarrow C.\textsc{get}(i,1)$
6:  while $(i+\textsc{ToString}(count)))$ is a key in $D$ do
7:   $count\leftarrow count+1$
8:  end while
9:  $key\leftarrow i+\textsc{ToString}(count)$
10:  $C[i]\leftarrow count+1$
11:  return keykey
12:end function
13:function DictAttack($Q_{m},k$)
14:  $W_{h}\leftarrow\textsc{Tokenize}(Q_{m})$
15:  $D\leftarrow\{\}$, $K_{h}\leftarrow[]$, $C\leftarrow\{\}$
16:  for each $w$ in $W_{h}$ do
17:   $key\leftarrow\textsc{GenerateUniqueKey}(w,D,C)$
18:   $D[key]\leftarrow w$
19:   $K_{h}.\textsc{append}(key)$
20:   $W_{nh}\leftarrow\textsc{GenerateHarmlessSynonyms}(w,k)$
21:   for each $s$ in $W_{nh}$ do
22:     $key_{nh}\leftarrow\textsc{GenerateUniqueKey}(s,D,C)$
23:     $D[key_{nh}]\leftarrow s$
24:   end for
25:  end for
[…]
27:  $P_{control}\leftarrow\textsc{FormatAsJSONSchema}(D)$
28:  return $(P_{data},P_{control})$
29:end function
The primary limitation of EnumAttack is that the malicious intent is explicitly visible in the grammar rules. To address this, we introduce DictAttack, a more sophisticated CDA that decouples the malicious payload across the data and control planes. DictAttack exploits the LLM’s ability to act as a reasoning engine that can map abstract keys to specific values provided in its context.
As detailed in Algorithm 1, DictAttack operates through a three-step generation process: (1) Tokenization and Obfuscation: The original malicious query $Q_{m}$ is tokenized into individual words $W_{h}$. (2) Dictionary Synthesis: A dictionary $D$ is constructed where each harmful word $w\in W_{h}$ is assigned a unique, benign key (e.g., a1, b2, c3). To further obscure the intent, we synthesize $k$ times more harmless synonyms $W_{nh}$ and include them in the dictionary. The resulting dictionary $D$ appears as a standard, benign technical vocabulary to any individual plane guardrail. (3) Dual-Plane Payload Generation:
  • The data-plane payload $P_{data}$ is a prompt containing only the benign keys (e.g., “a1+b2+c3”), instructing the model to translate these keys using the provided dictionary and execute the resulting command.
  • The control-plane payload $P_{control}$ is the JSON schema containing the dictionary $D$ embedded in its properties or descriptions.
During inference, the model receives the benign keys in the prompt and the dictionary in the grammar. The constrained decoding process then forces the model to select tokens that correspond to the values mapped to those keys in the dictionary. Neither the keys in the prompt nor the words in the dictionary are inherently harmful when viewed in isolation. Only when the model reconstructs the query by joining the dictionary values does the malicious intent emerge. This “Dual-Plane Decoupling” makes DictAttack exceptionally evasive, as a guardrail would need to semantically audit both planes simultaneously, potentially recalling all history context, to detect the threat. By increasing the synthesis ratio $k$, the attacker can dilute the ratio of harmful words in the dictionary, making detection via frequency analysis or semantic mismatch even more difficult.
Interleaved DictAttack. In practice, the attacker can further decouple DictAttack across time in multi-turn interactions. Specifically, the benign key sequence can be sent in an earlier turn (thus becoming part of the model’s conversational context / KV cache), while the dictionary-embedded schema is only provided in a subsequent request. This interleaving degrades many “dual-plane” auditing defenses that only inspect the current prompt+grammar pair, effectively reducing them back to single-plane checks. We refer to this variant as Interleaved DictAttack and evaluate its impact in §§§ 5.3.4.
4.3. Other CDAs
Beyond the attacks presented so far, other forms of CDAs are possible. For example, EnDec (enforcedecode2024acl) and JailMine (jailmine2024arxiv), though designed as white-box methods that directly manipulate output logits, can also be realized indirectly via guided grammar.
APT (APT2025arxiv) leverages GuidedRegex to iteratively block refusal tokens, but the approach is non-functional in practice since its claimed backends, xgrammar (xgrammar2024arxiv) and outlines (outlines2023arxiv), lack support for negative lookahead regex (e.g., “(?!Sorry)”). Even if supported, the method is inefficient, requiring exhaustive token-by-token trials.
Prompt-based jailbreaks remain orthogonal to our work and can be combined with CDAs. Template-based methods (MasterKey (masterkey2024ndss), LLMFuzzer (llmfuzzer2024usenixsecurity), PAIR (PAIR2024arxiv), TAP (TAP2024nips)) can strengthen EnumAttack by crafting more complex “yes-prefix” fields, while linguistic (drattack2024emnlpfinding; puzzler2024aclfinding) and encoding attacks (zulu2024arxiv; base64attack2023nips; wei2023jailbroken) can further obscure malicious content by exploiting the mismatch between models’ advanced capabilities and their safety alignment.
Except JSON Schema, another powerful structured output option is guided grammar, which is defined via Extended Backus-Naur Form (EBNF), a meta-syntax used to express context-free grammars, which is effectively unbounded. This flexibility can also enable adversaries to craft complex, seemingly harmless grammars that can steer LLM generation toward arbitrary malicious goals. By formalizing attacks as grammar rules, CDAs represent a shift from probabilistic prompt engineering to deterministic control-plane manipulation, highlighting the urgent need to systematically analyze constrained decoding mechanisms and develop mitigations that can audit the semantic implications of a grammar, as we discuss in § 7.
5. Evaluation
Table 1. Summary of Datasets used for jailbreak attack evaluation, $\circ$ denotes w/o such property.
Dataset Size Category Extra Attack
AdvBench(advbench2023emnlp) 520 $\circ$ $\circ$
StrongREJECT(strongreject2024iclrws) 311 6 $\circ$
JailbreakBench(jailbreakbench2024nips) 100 10 $\circ$
HarmBench(harmbench2024arxiv) 100 3 $\circ$
[…]
CDAs represent a shift in the jailbreaking paradigm by targeting the decoding process rather than relying on prompt engineering. By manipulating the structured output specifications through carefully crafted schema definitions (as demonstrated in our EnumAttack implementation), these attacks force models to generate content that would otherwise be filtered by both external guardrails and internal alignment mechanisms. We have revealed the vulnerability within the grammar control plane through the proposed attacks (EnumAttack and DictAttack), where plain jailbreak questions can be embedded in the grammar to cause jailbreaks, while no guardrails currently in this vulnerable path.
In this section, we first build a LLM guardrail system to mitigate CDAs, which is shown in Figure 6. Following MasterKey (masterkey2024ndss), which already establish a LLM system with prompt-based guardrail, we further enrich a grammar guardrail to protect the control plane as mitigation strategy. We provide a systematic evaluation of EnumAttack and DictAttack across different models, benchmarks and guardrail choices.
5.1. Experimental Setup
Attack methods. Besides our proposed EnumAttack (§§ 4.1) and DictAttack (§§ 4.2), we evaluate five baseline data-plane attacks: IJP (jailbreak2024ccs), DrAttack (drattack2024emnlpfinding), Puzzler (puzzler2024aclfinding), Zulu (zulu2024arxiv), and Base64 (base64attack2023nips). For comparative analysis against SOTA data-plane jailbreaks, we also include PAIR (PAIR2024arxiv), TAP (TAP2024nips), and AutoDAN-Turbo (autodanturbo2025iclr). For each benchmark, we follow the methodology described in §§ 4.1 and §§ 4.2, applying EnumAttack and DictAttack to the harmful query intents provided by the respective dataset.
Datasets. Following previous works  (align2023nips; GCG2023arxiv; enforcedecode2024acl; JBShield2025usenixsecurity), we evaluate the performance of constrained decoding based attacks (CDAs) in six well-known benchmarks, whose statistics are shown in Table 1.
System-level Safeguards. To evaluate the attack’s effectiveness against real-world deployed systems, we construct a multi-layered defense environment:
  • Industrial Guardrails: We deploy llama-guard-3-8b (llamaguard2023arxiv) and OpenAI Moderation API (openaimoderation2023aaai), which are widely used for filtering prompts and responses.
  • SOTA Academic Defenses: We evaluate against recent JBShield (JBShield2025usenixsecurity) and SelfDefend (selfdefend2025security), two state-of-the-art data-plane defenses proved to be effective on known jailbreak attacks.
  • Dual-Plane Guard: We adapt SelfDefend (selfdefend2025security) to jointly audit the prompt and grammar, instantiated with gpt-4o as the auditing LLM (i.e., 4o guarding 4o when evaluating gpt-4o).
Large language models. We evaluate a broad spectrum of 13 LLMs with native structured output support, categorized by their scale and deployment scenarios. For proprietary models, we test the widely-deployed gpt-4o, gpt-4o-mini, and gemini-2.0-flash, alongside the latest flagship models gpt-5 and gemini-2.5-pro. For open-weight models, we evaluate a range of architectures including llama-3.1-8b (llama3-2024website), qwen-2.5-32b (qwen2.5_2024arxiv), mistral nemo (mistralnemo2024web), phi-3.5-moe (phi3safety2024arxiv), and gemma-2-9b (gemma2_2024arxiv), as well as high-capability models like deepseek-v3/r1 (deepseekv3-2025arxiv; deepseekr1-2025arxiv) and gpt-oss-120b (gpt-oss). Throughout Evaluation, we report results by attack scenario: EnumAttack is primarily evaluated on mainstream deployed systems (gpt-4o, gpt-4o-mini, gemini-2.0-flash), while DictAttack is highlighted on flagship and high-capability models (gpt-5, gemini-2.5-pro, deepseek-r1, gpt-oss-120b). All models [^1] Checkpoints used: gpt-4o-2024-0806, gpt-4o-mini-2024-0718, gemini-2.0-flash-001, gpt-5, gemini-2.5-pro, deepseek-v3-0324, deepseek-r1-0528, and gpt-oss-120b. For all open-weight models, we specifically evaluate their instruction-tuned versions (e.g., llama-3.1-8b-instruct, qwen-2.5-32b-instruct, gemma-2-9b-it) to ensure standard safety alignment is active. We also exclude preview version LLMs due to their instability, like gemini-3.0-pro-preview. are evaluated using black-box API access to simulate realistic attack scenarios.
Evaluation metrics. We first leverage the mainstream metric: Attack Success Rate (ASR), which quantifies the percentage of successful jailbreaks. ASR is a commonly used measure in LLM security research (GCG2023arxiv; align2023nips; enforcedecode2024acl; strongreject2024iclrws), typically relying on an external LLM to judge. Following this convention, we employ another powerful LLM gpt-4o (gpt4o2024web) as the judge in our evaluation, which is in line with previous work (strongreject2024iclrws). An attack is deemed successful only if the attacker’s query doesn’t trigger any safety flag during prompt and grammar auditing (i.e., bypassing guardrails), and the LLM itself answers the query as expected (i.e., bypassing internal alignment). Furthermore, for the guardrails deployed to defend against our attacks, we report their Detection Rate. This metric reflects the guardrail’s ability to identify malicious inputs; therefore, a lower detection rate indicates a more effective attack that successfully evades the defense mechanism.
For local open-weight LLMs, we test both vllm 0.10.1 (vllm2023sosp) and sglang 0.4.10 (sglang2024arxiv) for serving. Both vllm and sglang support an OpenAI compatible server with structured output features supported by a grammar backend like xgrammar (xgrammar2024arxiv) or outlines (outlines2023arxiv). We use the OpenAI compatible server to call the local LLMs so that the evaluation for local models is consistent with the evaluation for proprietary models. For large-scale open-weight models such as deepseek-v3/r1 and gpt-oss-120b, due to resource limits, we utilize APIs provided by OpenRouter (openrouter) to evaluate them in realistic production serving scenarios.
5.2. CDAs Can Effectively Bypass Prompt Guardrails
In this subsection, we focus on evaluating the effectiveness of CDAs against prompt guardrails, thus only the prompt guard of Figure 6 is enabled. We evaluate along three dimensions:
  1. (i)
    EnumAttack on prompt-guarded models. We test open-weight and proprietary LLMs under a state-of-the-art prompt guard, JBShield (JBShield2025usenixsecurity), using AdvBench as the task suite; we report ASR and guard detection rates.
  2. (ii)
    Comparison to prompt-based jailbreaks. Using llama-3.1-8b, we reproduce five representative prompt attacks (jailbreak2024ccs; drattack2024emnlpfinding; puzzler2024aclfinding; zulu2024arxiv; base64attack2023nips) under the same detection setting as JBShield (JBShield2025usenixsecurity), contrasting control-plane (schema) vs. data-plane (prompt) vulnerabilities.
  3. (iii)
    Infrastructure sensitivity. We ablate serving stack and grammar backend, covering two serving frameworks (vLLM and SGLang) and two grammar engines ([xgrammar] and [outlines]). We measure whether CDA effectiveness is tied to a particular implementation.
5.2.1. EnumAttack on prompt-guarded models.
Table 2. EnumAttack ASR on open-weight and proprietary LLMs on AdvBench. Results w/ and w/o prompt guardrail are listed outside/inside parentheses accordingly.
Model ASR
w/ (w/o)
phi-3.5-moe (phi3safety2024arxiv) 99.2% (99.2%)
mistral nemo (mistralnemo2024web) 100.0% (100.0%)
qwen-2.5-32b (qwen2.5_2024arxiv) 99.0% (99.0%)
llama-3.1-8b (llama3-2024website) 99.2% (99.2%)
gemma-2-9b (gemma2_2024arxiv) 98.6% (98.6%)
deepseek-v3 (deepseekv3-2025arxiv) 98.7% (98.7%)
gpt-4o (gpt4o2024web) 100.0% (100.0%)
gpt-4o-mini (gpt4omini2024web) 99.0% (99.0%)
gemini-2.0-flash (gemini2023arxiv) 95.8% (95.8%)
As shown in Table 2, EnumAttack achieves near-100% ASRs across all evaluated open-weight and proprietary models protected by JBShield. This outcome underscores EnumAttack’s high effectiveness to break existing prompt-based guardrails.
5.2.2. Comparison to prompt-based jailbreaks.
Table 3. ASR(%) on llama-3.1-8b with JBShield (JBShield2025usenixsecurity). While other attacks (jailbreak2024ccs; drattack2024emnlpfinding; puzzler2024aclfinding; zulu2024arxiv; base64attack2023nips) are effectively mitigated, JBShield is fully bypassed by our EnumAttack, which achieves even higher ASR by embedding the malicious query in the JSON schema payload.
Method IJP(jailbreak2024ccs) DrAttack(drattack2024emnlpfinding) Puzzler(puzzler2024aclfinding) Zulu(zulu2024arxiv) Base64(base64attack2023nips)
Eval Size 820 820 20 820 490
NoDef 48.9% 65.9% 100% 2.7% 8.4%
JBShield 0.85% 0% 0% 0% 0.82%
Ours+NoDef 47.8% 74.1% 100% 66.8% 67.9%
Ours+JBShield 47.8% 74.1% 100% 66.8% 67.9%
As depicted in Table 3, JBShield detects prompt-based attacks with very high accuracy, leading to near 0% ASR regardless of the original attack’s performance, demonstrating effectiveness against traditional data-plane threats. However, JBShield fails to detect EnumAttack, achieving identical results with/without guard. This discrepancy arises because JBShield is a lightweight defense that relies on static attack profiles and model-specific white-box states of the prompt. By maintaining a completely benign prompt body, CDAs like EnumAttack completely bypass its detection mechanism, rendering such static, data-plane-only defenses inapplicable to control-plane vulnerabilities. Notably, the lengthy question body in IJP dataset (jailbreak2024ccs) doesn’t fit well in the template of EnumAttack, leading to 47.8% ASR in one query. While the judge model ft-mistral-sorrybench-2024-06 (sorrybench2025iclr) doesn’t know Zulu and Base64 languages, and the ASRs are therefore not accurately assessed, leading to lower ASR, which is an inherent limitation for the LLM-as-a-judge method during jailbreak evaluation. As a result, we use gpt-4o as a more powerful and accurate LLM judge for all other evaluations as stated in §§ 5.1.
5.2.3. Infrastructure sensitivity.
Figure 5. EnumAttack evaluation of using different grammar backends and serving engines with llama-3.1-8b, except negligible differences caused by temperature and minor variations in the backend, the attack is consistently successful.
To ensure our findings are not backend-specific, we evaluate two grammar engines (xgrammar (xgrammar2024arxiv) and outlines (outlines2023arxiv)) across mainstream serving stacks (vllm (vllm2023sosp) and sglang (sglang2024arxiv)). As shown in Figure 5, EnumAttack remains consistently successful across all combinations, demonstrating that the vulnerability is inherent to the mechanism rather than a specific implementation flaw.
Figure 6. Overview of our CDA mitigation system, where prompt guard, grammar guard and LLM internal alignment works together to mitigate various CDAs, like DictAttack presented here.
Table 4. Major ASR results on mainstream deployed models (gpt-4o, gpt-4o-mini, gemini-2.0-flash). Results w/ and w/o grammar guardrail are listed outside/inside parentheses accordingly. We use llama-guard-3-8b as the grammar guardrail and set k=1k{=}1 for DictAttack.
Model Method AdvBench HarmBench JailbreakBench SorryBench StrongREJECT
gpt-4o Baseline 1.2%(1.2%) 26.0%(26.0%) 10.0%(10.0%) 33.9%(33.9%) 5.1%(5.1%)
EnumAttack 3.4%(100.0%) 7.0%(100.0%) 8.0%(100.0%) 16.6%(96.4%) 5.1%(99.4%)
DictAttack 98.1%(99.8%) 98.3%(100.0%) 98.3%(100.0%) 96.9%(98.6%) 96.7%(98.4%)
gpt-4o-mini Baseline 2.1%(2.1%) 44.0%(44.0%) 14.0%(14.0%) 42.5%(42.5%) 8.9%(8.9%)
EnumAttack 3.4%(99.0%) 7.0%(100.0%) 8.0%(97.0%) 16.8%(96.4%) 5.1%(98.1%)
DictAttack 95.5%(97.1%) 83.6%(85.0%) 92.4%(94.0%) 83.8%(85.2%) 69.4%(70.6%)
gemini-2.0-flash Baseline 17.5%(17.5%) 46.0%(46.0%) 22.0%(22.0%) 29.3%(29.3%) 6.0%(6.0%)
EnumAttack 3.3%(95.8%) 7.0%(92.0%) 8.0%(92.0%) 15.5%(89.5%) 4.2%(81.2%)
DictAttack 96.8%(98.5%) 97.3%(99.0%) 89.5%(91.0%) 93.9%(95.5%) 91.4%(93.0%)
Table 5. DictAttack ASR on flagship models (gpt-5, gemini-2.5-pro, deepseek-r1, gpt-oss-120b). Results w/ and w/o grammar guardrail are listed outside/inside parentheses accordingly.
Model Adv Harm JB Sorry SR
gpt-5 94.4%(96.0%) 90.4%(92.0%) 93.4%(95.0%) 91.6%(93.2%) 93.9%(95.5%)
gemini-2.5-pro 97.5%(99.2%) 98.3%(100.0%) 98.3%(100.0%) 97.6%(99.3%) 97.3%(99.0%)
deepseek-r1 97.2%(98.9%) 93.4%(95.0%) 95.4%(97.0%) 96.9%(98.6%) 96.4%(98.1%)
gpt-oss-120b 96.6%(98.3%) 90.4%(92.0%) 89.4%(91.0%) 94.8%(96.4%) 94.2%(95.8%)
5.3. Evaluation against Grammar-guarded Systems
In this section, we systematically evaluate the effectiveness of CDAs against grammar-based guardrails across two dimensions:
  1. (i)
    Attack effectiveness under grammar-level auditing. We test EnumAttack and DictAttack across a comprehensive set of models, including mainstream proprietary models (gpt-4o, 4o-mini, gemini-2.0-flash), recent flagship models gpt-5 and gemini-2.5-pro, and high-capability open-source models deepseek-r1 and gpt-oss-120b.
  2. (ii)
    DictAttack with combined prompt and grammar guarding. We further evaluate the evasiveness of DictAttack against a strengthened Dual-Plane Guard to assess whether cross-plane security coordination can mitigate the threat.
5.3.1. Comparative Evasiveness: CDA vs. SOTA Data-Plane Attacks
To further validate the robustness of our selected guardrails and highlight the unique threat posed by CDAs, we benchmark the fundamental evasiveness of DictAttack against well-known SOTA data-plane jailbreaks: PAIR (PAIR2024arxiv), TAP (TAP2024nips), and AutoDAN-Turbo (autodanturbo2025iclr).
We conducted this evaluation on gpt-4o using the AdvBench dataset under two settings:
  • Undefended: Direct attack against the model’s internal alignment.
  • Guarded: Data-plane attacks use the best baseline guard; DictAttack is evaluated with our Dual-Plane Guard enabled.
Table 6. Comparison of Attack Success Rate (ASR) between SOTA data-plane attacks and DictAttack on gpt-4o. While standard guardrails effectively mitigate powerful traditional attacks, DictAttack remains highly evasive.
Attack Method ASR (Undefended) ASR (Guarded)
PAIR 11.5% 1.6%
TAP 24.4% 0.4%
AutoDAN-Turbo 96.0% 26.0%
[…]
DictAttack is evaluated with our Dual-Plane Guard (SelfDefend-based; gpt-4o auditor); others use the best data-plane guard.
The results in Table 6 reveal a critical disparity. Powerful traditional attacks like AutoDAN-Turbo, while highly effective against undefended models (96.0In sharp contrast, DictAttack not only achieves near-perfect success undefended (99.8This underscores that while our system guardrails are effective against known data-plane threats, they are fundamentally insufficient against the cross-plane decoupling of DictAttack. A more detailed analysis of DictAttack’s performance across different configurations is provided in §§§ 5.3.2.
5.3.2. EnumAttack and DictAttack on grammar-guarded models
Table 4 and Table 5 report ASRs across five benchmarks. The baseline uses each benchmark’s original prompts, while EnumAttack and DictAttack wrap the same intents using our schema-based approach. For each entry, the value outside parentheses is the ASR with the grammar guard enabled and the value in parentheses is the ASR without the grammar guard.
First, baseline results confirm that direct prompting is largely ineffective against the safety alignment of modern LLMs. Although different benchmarks cover different aspects of harmful content, the overall baseline ASR remains consistently low, with averages ranging from 6.7% to 38.7% across datasets. These measurements validate the effectiveness of conventional safety alignment against plain jailbreak attacks.
Second, while EnumAttack achieves an average ASR of 95.8% undefended, simple grammar auditing reduces its success rate to 7.9% (Table 4). This confirms that its reliance on explicit literals makes it highly detectable, which motivates the need for more evasive CDAs like DictAttack.
Third, DictAttack remains highly effective even under grammar auditing. The same grammar-level defense is almost entirely ineffective against DictAttack. As shown in Table 4 and Table 5, DictAttack maintains high ASRs across both mainstream deployed models and flagship/high-capability models. For example, averaged across the five benchmarks, DictAttack achieves 94.3–99.5% ASR on gpt-5, gemini-2.5-pro, deepseek-r1, and gpt-oss-120b without grammar guard (inside parentheses); applying llama-guard-3-8b with k=1k{=}1 (1.7% detection rate) yields 92.7–97.8% ASR with guard (outside parentheses). This demonstrates that the vulnerability is not a flaw of older checkpoints but a fundamental issue that persists in the latest, most heavily aligned models. The slightly lower ASR for gpt-4o-mini (86.4%) is attributed to its reduced reasoning capability, which sometimes fails to correctly reconstruct the query from the dictionary rather than any safety-related refusal.
5.3.3. DictAttack with combined prompt- and grammar-guarding
The failure of the grammar-only guard is straightforward: DictAttack decouples the malicious query into two benign-looking components—one in the data plane (prompt) and the other in the control plane (grammar). Individual guardrails, which only examine one plane at a time, cannot identify the combined threat. Theoretically, only by auditing both planes together can a guard potentially detect the jailbreak intent.
Thus, we further evaluate a Dual-Plane Guard setting, adapted from the SOTA SelfDefend (selfdefend2025security). We adapt SelfDefend’s audit prompt to take both the data-plane prompt (keys) and the control-plane grammar (dictionary), using gpt-4o as the auditing LLM.
In this subsection, we evaluate DictAttack across synthesis ratios $k$ using multiple moderation models under two defense policies: Defense 1 (grammar-only) and Defense 2 (combined prompt- and grammar-guarding). Table 7 reports detection rates across configurations; lower detection rates indicate reduced guard effectiveness.
First, the results demonstrate that even a unified audit fails to effectively mitigate DictAttack. While llamaguard can detect 28.6% of attacks, Moderation APIs almost detect none, probably due to restricted model capabilities. More strikingly, the state-of-the-art SelfDefend (selfdefend2025security)—even with its best strategy—only achieves an 18.5% detection rate against DictAttack on AdvBench. This failure occurs because, while an LLM judge is theoretically capable of reasoning, modern structured output schemas are highly complex and difficult for even powerful LLMs to “parse” and audit semantically in a unified way. The spatial separation of the payload across planes creates a reasoning gap that current auditing-based defenses cannot reliably bridge.
Second, increasing the synthesis factor $k$ markedly weakens the detection. Under Defense 1, gpt-4o attains $\approx 18\%$ detection at $k=0$; however, when $k$ is raised to 10, detection drops to 5.3%, as the malicious tokens are further diluted within the benign dictionary. This reveals that DictAttack exploits a fundamental blind spot: the lack of cross-plane coordination in LLM safety architectures.
Table 7. Detection rate on AdvBench (advbench2023emnlp) under Defense 1 (grammar-only) and Defense 2 (Grammar + Prompt Dual Guard, adapted from SelfDefend; gpt-4o auditor).
Defense 1: Grammar Guard
Attack Method llamaguard Moderation gpt-4o
EnumAttack 97.7% 90.2% 100.0%
DictAttack ($k=0$) 6.5% 0.0% 17.7%
DictAttack ($k=1$) 1.7% 0.0% 8.1%
DictAttack ($k=10$) 1.5% 0.0% 5.3%
Defense 2: Grammar + Prompt Dual Guard
Attack Method llamaguard Moderation gpt-4o
DictAttack 28.6% 0.2% 24.2%
Interleaved DictAttack 1.5% 0.0% 5.3%
5.3.4. Discussion: Space-Time Decoupling via Interleaved DictAttack
While the Dual-Plane Guard theoretically attempts to bridge the reasoning gap, its sensitivity remains low (24.2% detection rate for gpt-4o). This is because, in addition to the spatial decoupling between data and control planes, an adversary can further exploit the temporal dimension via Interleaved DictAttack.
In this variant, the attacker splits the two DictAttack payloads across turns: the prompt payload (keys) is sent first and stored in the conversation context (KV cache), while the grammar payload (dictionary) arrives later in a separate request. Because practical dual-plane guards are typically implemented as per-request auditors, this temporal decoupling effectively reduces a dual-plane audit to a single-plane check, yielding only 5.3% detection for gpt-4o. Mitigating this would require cross-request retrieval and auditing of the entire conversation history at every turn, which is impractical:
  • Unacceptable Cost: Auditing the full history at every turn would require tremendous computational overhead, as the guard system doesn’t know a priori when a malicious payload will arrive.
  • Reasoning Complexity: As the conversation grows (e.g., up to gpt-4o’s 128k token limit), the guard model would require context length and reasoning capabilities comparable to or greater than the protected model itself, which is prohibitively expensive for a real-time monitor.
  • Interpretability: Modern JSON schemas are often dense and opaque. Expecting a guard LLM to semantically reconstruct a query from a dictionary buried 50 turns back in history represents a “needle in a haystack” problem that current systems are not designed to solve.
In conclusion, DictAttack demonstrates that by decoupling intent across both space (planes) and time (requests), adversaries can render external auditing defenses heavy, expensive, and ultimately impractical.
5.4. Case study: how internal alignment is broken by malicious grammar
Figure 7. Token distribution case study ablating progressive attack methods, the exact probability distribution is sampled from phi-3.5-moe model with a sequence_length = 5. Refusal tokens, safe tokens and jailbreak tokens are marked explicitly with their generation probability, less-than-1e-16 values are omitted.
Figure 8. Quantitative evaluation of token distribution shift with progressive attack methods. With more methods, the token distribution shifts from a safety-aligned one (with significantly high refusal token probability dominating the answer) to a balanced, context-sensitive one (diverse token choices answering the question, where no token choice is dominating as measured by token probability value).
CDA vulnerabilities are two-fold: (1) External Guardrail failure (bypassing filters) and (2) Internal Alignment failure (coercing models to generate harmful content instead of refusing). To analyze the latter, we evaluate token probability distributions using phi-3.5-moe (phi3safety2024arxiv) across three categories: refusal, safe, and jailbreak tokens.
Figure 7 illustrates a dramatic logit shift toward jailbreaking across progressive methods:
  1. (1)
    Direct Prompt: The model refuses with near 100% probability.
  2. (2)
    Yes-prefix: Minimal impact; the model still strongly resists.
  3. (3)
    JSON format: Refusal probability drops by approximately half.
  4. (4)
    EnumAttack: Direct refusals vanish, but safe tokens remain highly probable.
  5. (5)
    System Prompt: Combined with EnumAttack, this suppresses safe tokens, making jailbreak tokens the most likely outcome.
  6. (6)
    Chain EnumAttack: Using malicious context from a weaker model, even this strongly aligned model is fully jailbroken; the top-5 beams are all jailbreak tokens.
  7. (7)
    Chained Steps: With deeper chained steps, the alignment is broken completely.
Quantitatively (Figure 8), token proportions shift from refusal to safe, and finally to jailbreak tokens. This transition reveals how CDAs manipulate the model’s internal probability space. While pre-trained LLMs naturally exhibit a diverse and open-ended token distribution, safety alignment artificially distorts this space by forcing refusal tokens to dominate when harmful intents are detected. By imposing structured constraints that fall outside the model’s safety-aligned training distribution, CDAs allow the generation process to “escape” these artificial refusal peaks. Consequently, the distribution reverts from a safety-aligned state (dominated by refusal tokens, ¿90%) to a balanced, context-sensitive state, restoring the model’s original open-ended generation capacity and enabling the production of harmful content with high probability.
This observation strongly echoes the findings of deepalignment2025iclr, which argue that current LLM safety alignment is often “only a few tokens deep.” By leveraging grammar constraints to force these critical first few tokens—such as affirmative prefixes or malicious intent literals—CDAs effectively bypass this shallow layer of internal resistance. Once the initial safety-critical token selection is overridden via deterministic control-plane manipulation, the model’s subsequent generation behaves as if no alignment were present, highlighting a fundamental fragility in current safety-training paradigms that rely primarily on steering the start of the response.
Table 8. Summary of existing jailbreak attacks adapted from  (JBShield2025usenixsecurity), - indicates the method does not use the listed resource or lacks that capability, $\circ$ denotes white-box attack and $\bullet$ denotes black-box attack.
Categories Jailbreaks Extra Assist White/Black box Target LLM Queries I/O-Based
Manually-designed IJP(jailbreak2024ccs) Human $\bullet$ - Input
[…]
Template-based MasterKey(masterkey2024ndss) LLM $\bullet$ $\sim$200 Input
LLMFuzzer(llmfuzzer2024usenixsecurity) LLM $\bullet$ $\sim$500 Input
AutoDAN(autodan2024iclr) LLM $\circ$ $\sim$200 Input
PAIR(PAIR2024arxiv) LLM $\bullet$ $\sim$20 Input
TAP(TAP2024nips) LLM $\bullet$ $\sim$20 Input
StructTransform(structtransform2025arxiv) LLM $\bullet$ $\sim$3 Input
[…]
Encoding-based Zulu(zulu2024arxiv) - $\bullet$ - Input
Base64(base64attack2023nips) - $\bullet$ - Input
[…]
6. Related Work
6.1. Jailbreak Attacks on LLMs
Jailbreak attacks aim to craft malicious inputs that cause LLMs to violate their safety guidelines. Following early suggestions by Carlini et al. (align2023nips), numerous jailbreak methods have emerged. We adopt the taxonomy from JBShield (JBShield2025usenixsecurity), which categorizes attacks as manual-designed, optimization-based, template-based, linguistics-based, and encoding-based. As shown in Table 8, we extend this taxonomy with a new category, output-based attacks, which is highly relevant to our work.
Manual-designed Jailbreaks involve manually crafting malicious inputs. Notable examples include the In-the-wild Jailbreak Prompts (IJP) (jailbreak2024ccs), which document real-world attempts shared on social media.
Optimization-based Jailbreaks use automated algorithms to craft adversarial prompts. GCG (GCG2023arxiv) adds an adversarial suffix to prompts using a greedy algorithm, while SAA (SAA2025iclr) combines templates with a random search strategy. While automated, these attacks typically require white-box or logit access and a large number of queries.
Template-based Jailbreaks embed harmful requests within sophisticated, optimized templates. Methods like MasterKey (masterkey2024ndss), LLMFuzzer (llmfuzzer2024usenixsecurity), AutoDAN (autodan2024iclr), PAIR (PAIR2024arxiv), and TAP (TAP2024nips) often use other LLMs to generate or refine these templates. Concurrent work like StructTransform (structtransform2025arxiv) also targets structured generation, aligning with our findings.
Linguistics-based Jailbreaks conceal malicious intent within seemingly benign inputs using linguistic properties. For example, DrAttack (drattack2024emnlpfinding) decomposes and reconstructs prompts, while Puzzler (puzzler2024aclfinding) uses combinations of diverse clues to bypass safety mechanisms.
Encoding-based Jailbreaks obfuscate malicious prompts by translating them into less common languages (Zulu (zulu2024arxiv)) or encoding formats like Base64 (base64attack2023nips), exploiting the mismatch between the model’s high-level capabilities and its restricted safety alignment as identified by foundational work (wei2023jailbroken).
Output-based Jailbreaks is an emerging category. White-box methods like EnDec (enforcedecode2024acl) and JailMine (jailmine2024arxiv) directly manipulate model logits or token probabilities to construct harmful outputs. Closely related to our work, APT (APT2025arxiv) uses the GuidedRegex feature in structured output APIs to build a prefix tree that explicitly bans refusal tokens, forcing a harmful response.
7. Discussions
Given the significant vulnerabilities revealed by CDAs, we discuss the implications for the LLM safety landscape and propose potential mitigations. As shown in Figure 6, current auditing practices focus on two phases: input and output auditing (outauditing2024arxiv).
Input Auditing. Input-focused strategies use classifiers or small LLMs to filter prompts in parallel with generation, providing a cost-effective defense. However, CDAs easily bypass this by hiding malicious intent within the control-plane grammar while keeping the data-plane prompt benign. While auditing the grammar alongside the prompt (Dual-Plane Guard) could mitigate this, it remains impractical for complex, interleaved multi-turn attacks due to high reasoning costs and the limited capability of lightweight guard models.
Figure 9. BenignEnumAttack (red) and a benign redteaming Q-A auditing process (blue). Existing output auditing methods cannot identify their differences, so that they will be refused or returned together, causing either false positive or false negative.
Output Auditing. Output auditing filters generated content but faces steep challenges in latency, cost, and false positives (youcant2025www). Consequently, major APIs like openai and gemini often omit it. Beyond practical limits, we question if output auditing is a “silver bullet”. To test this, we developed BenignEnumAttack (Figure 9), which mimics a benign structured output (e.g., a StrongREJECT report) while embedding harmful content. This creates a semantic deadlock: since the jailbreak and a legitimate safety audit appear structurally identical, an auditor must either trigger a false positive (refusing safety research) or a false negative (leaking the jailbreak). This indicates that output-based defense has fundamental conceptual limitations beyond mere implementation costs.
Mitigation Strategies. To secure the control plane, we propose several strategies across different layers:
  1. (1)
    Lightweight Heuristics: Filters could flag suspicious Literal-to-Logic Ratios or Semantic Mismatches between prompts and schemas. However, such heuristics face similar false positive risk in Figure 9. For example, a legitimate tool for generating a technical glossary or a translation table would naturally contain many string literals and might trigger these filters.Ensuring safety without breaking legitimate utility remains a complex trade-off that requires more robust and context-aware auditing.
  2. (2)
    Grammar-Plane Auditing: Direct auditing of JSON Schemas can mitigate basic attacks like EnumAttack. However, cross-plane coordination for DictAttack remains a high-cost reasoning challenge.
  3. (3)
    Context-Aware Token Attribution: Current output auditing processes cannot distinguish between user-prefilled content and model-generated content—a vulnerability motivated by BenignEnumAttack. A more robust approach would implement token-level provenance tracking during the generation process, enabling auditing systems to differentiate between tokens originating from enum constraints (or others specified by grammars) versus those freely generated by the model. This would enable more accurate detection of attempts to manipulate the generation trajectory through constrained fields, neutralizing threats like BenignEnumAttack.
  4. (4)
    Safety-Preserving Constraints: We suggest whitelisting refusal tokens (e.g., “sorry”) to override user-defined grammars, ensuring internal safety alignment can trigger even within constrained outputs when jailbreak is absolutely detected. However, implementation challenges include balancing strict grammar adherence with safety requirements, as excessive safety tokens might disrupt valid structured outputs in certain contexts.
  5. (5)
    Integrated Safety Signaling: Co-design LLM with auditing signal tokens: we still whitelist some tokens in the constrained decoding process, but we whitelist special tokens that indicate potential auditing signals. For example, we can co-design LLMs to generate special auditing tokens (e.g., <unsafe>) when encountering harmful contexts, to enable real-time, constraint-agnostic monitoring.
8. Conclusion
In this study, we introduced the Constrained Decoding Attack (CDA), revealing a critical control-plane vulnerability in LLMs. Our evaluation across 13 models, including gpt-5 and gemini-2.5-pro, demonstrates that by weaponizing deterministic grammar constraints, attackers can bypass both internal alignment and external guardrails with near-perfect success. More importantly, our primary contribution, DictAttack, exposes a fundamental “semantic gap” in current defenses; by decoupling intent across space and time, it maintains a 75.8% ASR even against state-of-the-art jailbreak guardrails. By revealing this previously unexplored attack surface, our work contributes to developing more comprehensive security paradigms for LLMs that address safety at all stages of LLM generation.
Ethical Considerations
Our work identifies significant security vulnerabilities in constrained decoding mechanisms for LLMs by demonstrating how guided grammar manipulation can bypass safety guardrails. We empirically show that state-of-the-art models remain vulnerable to Constrained Decoding Attacks despite robust safety alignment efforts. The security implications are particularly concerning as CDAs exploit a fundamental gap in existing safety architectures, which primarily focus on input filtering rather than the generation process itself.
By publishing this research, we aim to enable the security community to address these vulnerabilities before they can be widely exploited. Since CDAs target low-level decoding processes exposed through structured output features, mitigating these attacks presents unique challenges distinct from traditional jailbreak defenses. We have acknowledged our findings to affected model providers like OpenAI and Gemini and propose several mitigation strategies that could reduce risk without compromising legitimate structured output functionality. We believe our analysis provides crucial insights for developing more comprehensive safety mechanisms that protect the entire generation pipeline rather than just input and output boundaries.

### v3

@@
-Beyond Prompts: Space–Time Decoupling Control-Plane Jailbreaks in LLM Structured Output
+When Grammar Guides the Attack: Uncovering Control-Plane Vulnerabilities in LLMs with Structured Output
 Abstract.
 Content Warning: This paper may contain unsafe or harmful content generated by LLMs that may be offensive to readers.
-Large Language Models (LLMs) are extensively used as tooling platforms through structured output APIs to ensure syntax compliance so that robust integration with existing software, like agent systems, can be achieved.
-However, the feature enabling the functionality of grammar-guided structured output presents significant security vulnerabilities.
-In this work, we reveal a critical control-plane attack surface orthogonal to traditional data-plane vulnerabilities.
-We introduce Constrained Decoding Attack (CDA), a novel jailbreak class that weaponizes structured output constraints to bypass both external auditing and internal safety alignment.
-Unlike prior attacks focused on input prompt designs, CDA operates by embedding malicious intent in schema-level grammar rules (control-plane) while maintaining benign surface prompts (data-plane).
-We instantiate this with two proof-of-concept attacks: EnumAttack, which embeds malicious content in enum fields; and the more evasive DictAttack, which decouples the malicious payload across a benign prompt and a dictionary-based grammar.
-Our evaluation spans a broad spectrum of 13 proprietary/open-weight models.
-In particular, DictAttack achieves 94.3–99.5% ASR across five benchmarks on gpt-5, gemini-2.5-pro, deepseek-r1, and gpt-oss-120b.
-Furthermore, we demonstrate the significant challenge in defending against these threats: while basic grammar auditing mitigates EnumAttack, the more sophisticated DictAttack maintains a 75.8% ASR even against multiple state-of-the-art jailbreak guardrails.
-This exposes a critical “semantic gap” in current safety architectures and underscores the urgent need for cross-plane defenses that can bridge the data and control planes to secure the LLM generation pipeline.
+Large Language Models (LLMs) increasingly serve as tooling platforms through structured output APIs, but the grammar-guided decoding that powers this feature opens a critical control-plane attack surface orthogonal to traditional data-plane vulnerabilities.
+We introduce Constrained Decoding Attack (CDA), a new jailbreak class that targets the LLM control plane.
+CDA is best characterized as a control-to-semantic pipeline: (1) schema-enforced logit masking injects a malicious prefix into the generation trajectory, and (2) the model itself completes the harmful intent.
+Unlike data-plane jailbreaks that rely on bypassing alignment with visible inputs, CDA acts on the decoding process itself, so internal safety alignment alone cannot stop it.
+We instantiate CDA with EnumAttack, which hides malicious content in enum fields, and the more evasive DictAttack, which decouples the payload across a benign prompt and a dictionary-based grammar.
+Across 13 proprietary/open-weight models and five standard benchmarks, DictAttack achieves 94.3–99.5% Attack Success Rate (ASR) on flagship models including gpt-5, gemini-2.5-pro, deepseek-r1, and gpt-oss-120b.
+While basic grammar auditing mitigates EnumAttack, DictAttack still sustains 75.8% ASR against SOTA jailbreak guardrails, exposing a “semantic gap” that demands cross-plane defenses bridging the data and control planes.
+Project page and code are available here.
 large language models, structured output, constrained decoding, jailbreak, guardrails, security
+conference: The 2026 ACM SIGSAC Conference on Computer and Communications Security; November 15-19, 2026; Hague, Netherlandssubmissionid: 752ccs: Security and privacy Social aspects of security and privacyccs: Computing methodologies Natural language generation
 1. Introduction
-Large language models (LLMs) (gpt4o2024web; gpt4techreport2023arxiv; claude2024website; gemini2023arxiv; llama2023arxiv; llama2-2023arxiv; llama3-2024website; qwen2.5_2024arxiv; phi3safety2024arxiv) have demonstrated remarkable capabilities across a wide range of tasks, from coding assistance to open-domain dialogue.
+Large language models (LLMs) (OpenAI, 2024a; Achiam et al., 2023; Anthropic, 2024; Gemini Team Google, 2023; AI, 2023b; Meta AI, 2023, 2024; Qwen, 2025; Haider et al., 2024) have demonstrated remarkable capabilities across a wide range of tasks, from coding assistance to open-domain dialogue.
 Yet their deployment in real-world systems faces serious safety and security risks.
 Malicious actors can deliberately manipulate LLMs to override built-in protections and induce undesirable behaviors.
-Such attempts, commonly known as jailbreaks(harmbench2024arxiv; sorrybench2025iclr; promptengineering2023arxiv; align2023nips; GCG2023arxiv; autodan2024iclr; SAA2025iclr), can produce harmful outputs such as misinformation, phishing content, or hate speech.
-Attackers achieve this through adversarial prompts (visualizejailbreak2023arxiv; align2023nips; GCG2023arxiv; jailbreak2024ccs), code injection (codeattack2024aclfinding), linguistic/ciphered inputs (base64attack2023nips; wei2023jailbroken), or manipulation of decoding parameters (exloitparam2024iclr).
-To mitigate these threats, AI (jailbreaksurvey2024aclfindings; safedecoding2024acl; enforcedecode2024acl; drattack2024emnlpfinding; SAA2025iclr; deepalignment2025iclr) and security researchers (jailbreak2024ccs; llmfuzzer2024usenixsecurity; JBShield2025usenixsecurity; masterkey2024ndss) have pursued two main defense strategies: internal safety alignment, which integrates guardrails directly into the model (phi3safety2024arxiv; deepalignment2025iclr; ji2023beavertails; pkusaferlhf; zhao2025improving), and external guardrails, which monitor and filter inputs or outputs in real-time (gpt3contentmoderation2023ijcai; llamaguard2023arxiv; contentmoderation2024ccs; safedecoding2024acl).
+Such attempts, commonly known as jailbreaks(Mazeika et al., 2024; Xie et al., 2025; Liu et al., 2024b; Carlini et al., 2023; Zou et al., 2023; Liu et al., 2024a; Andriushchenko et al., 2025), can produce harmful outputs such as misinformation, phishing content, or hate speech.
+Attackers achieve this through adversarial prompts (Qi et al., 2024; Carlini et al., 2023; Zou et al., 2023; Shen et al., 2024), code injection (Ren et al., 2024), linguistic/ciphered inputs (Wei et al., 2023a, b), or manipulation of decoding parameters (Huang et al., 2024).
+To mitigate these threats, AI (Xu et al., 2024b, a; Zhang et al., 2024; Li et al., 2024a; Andriushchenko et al., 2025; Qi et al., 2025) and security researchers (Shen et al., 2024; Yu et al., 2024; Zhang et al., 2025; Deng et al., 2024) have pursued two main defense strategies: internal safety alignment, which integrates guardrails directly into the model (Haider et al., 2024; Qi et al., 2025; Ji et al., 2023, 2024; Zhao et al., 2025), and external guardrails, which monitor and filter inputs or outputs in real-time (Wang et al., 2023; AI, 2023a; Wu et al., 2024; Xu et al., 2024a).
 Figure 1.
-(1) Prompt-based data plane jailbreak attack, mitigated by a guardrail, (2) EnumAttack, using structured output (LLM control plane) to embed malicious question, currently not guarded, (3) DictAttack, decoupling malicious payload into benign prompt and grammar, therefore jailbreaking system with both plane guardrails
+(1) Prompt-based data plane jailbreak attack, mitigated by a guardrail, (2) EnumAttack, using structured output (LLM control plane) to embed malicious question, currently not guarded, (3) DictAttack, decoupling malicious payload into benign prompt and grammar, therefore jailbreaking system with both plane guardrails.
 While existing defenses aim to safeguard model interactions at the prompt and output level, the paradigm of LLM deployment is shifting.
-Beyond standalone chatbots, LLMs are increasingly integrated into complex agent frameworks (langchain; langgraph) and enterprise automation pipelines (mcp2025website; cursor_ide) where they function as core reasoning engines.
-In these agentic workflows (e.g., Cursor (cursor_ide), LangChain (langchain), and Model Context Protocol (MCP) (mcp2025website)), a service component or an agent backend often supplies a grammar—typically a JSON schema—to a hosted LLM to ensure structured and reliable outputs.
+Beyond standalone chatbots, LLMs are increasingly integrated into complex agent frameworks (Chase, 2022; The LangChain Team, 2023) and enterprise automation pipelines (Team, 2025; Anysphere, Inc., 2023) where they function as core reasoning engines.
+In these agentic workflows (e.g., Cursor (Anysphere, Inc., 2023), LangChain (Chase, 2022), and Model Context Protocol (MCP) (Team, 2025)), a service component or an agent backend often supplies a grammar - typically a JSON schema - to a hosted LLM to ensure structured and reliable outputs.
 This functionality is implemented through two complementary approaches.
-On the model side via post-training on structured data (learnschema_2025acl), and on the infrastructure side via constrained decoding techniques applied during inference (outlines2023arxiv; xgrammar2024arxiv).
-Model-side approaches alone, however, face a key limitation: even strong LLMs can hallucinate or misinterpret complex formats without explicit constraints, which makes them unreliable for advanced tooling use cases (structuredoutputhallucination2024naacl; structuredoutputeval2025arxiv).
+On the model side via post-training on structured data (Lu et al., 2025), and on the infrastructure side via constrained decoding techniques applied during inference (Willard and Louf, 2023; Dong et al., 2025).
+Model-side approaches alone, however, face a key limitation: even strong LLMs can hallucinate or misinterpret complex formats without explicit constraints, which makes them unreliable for advanced tooling use cases (Ayala and Bechard, 2024; Geng et al., 2025).
 In contrast, constrained decoding (as shown in Figure 2) works by exposing a grammar that defines the expected output format and converting it into a context-free grammar.
 This ensures that every output strictly follows the required structure.
-Because of its reliability and broad applicability, constrained decoding has become the de facto standard, adopted by both proprietary LLM providers (structuredoutputexplanation2024openai) and open-source infrastructures (vllm2023sosp; sglang2024arxiv; xgrammar2024arxiv).
+Because of its reliability and broad applicability, constrained decoding has become the de facto standard, adopted by both proprietary LLM providers (OpenAI, 2024b) and open-source infrastructures (Kwon et al., 2023; Zheng et al., 2024; Dong et al., 2025).
 This reliance on constrained decoding, however, exposes a new and critical vulnerability.
-In this work, we introduce the Constrained Decoding Attack (CDA), a novel class of jailbreaks that targets the LLM control plane—the grammar that dictates the output structure.
-Unlike conventional jailbreak attacks on the data plane, hiding unsafe instructions inside otherwise benign prompts, CDA embeds malicious intent directly into the grammar itself.
-For example, with CDA, an attacker could design a grammar that always requires a harmful API call like delete_all_files() as part of the output.
-Even if the model refuses unsafe prompts at the data plane, the grammar forces it to produce the malicious call.
+This paper demonstrates the Constrained Decoding Attack (CDA), a new class of jailbreaks targeting the LLM control plane, i.e., the grammar that dictates the output structure.
+CDA is best characterized as a control-to-semantic attack pipeline with two stages: (i) control-plane injection, where schema-enforced logit masks force the model into a poisoned generative trajectory, and (ii) model-driven semantic continuation, where the model itself produces coherent harmful content from that trajectory.
+Unlike conventional jailbreak attacks on the data plane, hiding unsafe instructions inside otherwise benign prompts and relying on the model voluntarily producing harmful output despite alignment, CDA embeds malicious intent directly into the grammar itself.
+This deterministically shapes the output trajectory — so internal safety alignment alone cannot mitigate it once the control plane is reachable.
+For example, with CDA an attacker can hide a harmful request inside the JSON Schema and ask the model to “answer it step by step”; constrained decoding commits the request to the model’s context, and the model itself fluently produces operational instructions in the subsequent unconstrained fields, even when the user prompt stays benign.
 Because existing defenses focus almost exclusively on the data plane, this control plane attack surface remains largely unguarded, allowing CDA to bypass both internal alignments and external guardrails.
@@
 EnumAttack targets the enum property in JSON Schema to force malicious strings into the LLM’s generation context.
-While independent public discovery (EnumAttackBlog) have also noticed the potential for exploiting enum fields, we provide the first systematic characterization and implementation of this vulnerability.
+While independent public discovery (Priyanshu, 2024) have also discussed the potential for exploiting enum fields, we provide the first systematic characterization and implementation of this vulnerability.
 While effective against prompt-based guards, its reliance on string literals makes it susceptible to basic grammar auditing.
-To overcome this, we propose DictAttack, our primary contribution.
+We then demonstrate how DictAttack - the primary contribution of this work - can overcome the limitation of EnumAttack.
 DictAttack decouples the malicious payload across both the data and control planes.
@@
 The benign data-plane prompt then provides a sequence of keys that instructs the model to retrieve and assemble the hidden malicious query from the grammar-provided dictionary during decoding.
-By splitting the intent this way, DictAttack renders individual plane-level guardrails—and even many combined ones—completely ineffective.
-Our extensive evaluation across 13 state-of-the-art models confirms the devastating effectiveness of these attacks across five standard benchmarks (advbench2023emnlp; harmbench2024arxiv; jailbreakbench2024nips; sorrybench2025iclr; strongreject2024iclrws).
+By splitting the intent this way, DictAttack renders individual plane-level guardrails, and even many combined ones, completely ineffective.
+Our extensive evaluation across 13 state-of-the-art models confirms the devastating effectiveness of these attacks across five standard benchmarks (Chen et al., 2022; Mazeika et al., 2024; Chao et al., 2024a; Xie et al., 2025; Souly et al., 2024).
 In particular, under DictAttack, we observe 94.3–99.5% ASR across these benchmarks on flagship proprietary and open-weight models (gpt-5, gemini-2.5-pro, deepseek-r1, gpt-oss-120b).
 More importantly, we show that while standard grammar-auditing can mitigate the straightforward EnumAttack, it is largely bypassed by the more sophisticated DictAttack.
-We benchmark DictAttack against state-of-the-art industrial and academic guardrails (llamaguard2023arxiv; openaimoderation2023aaai; JBShield2025usenixsecurity; selfdefend2025security); even with a coordinated Dual-Plane Guard auditing both planes together, DictAttack maintains a high ASR of 75.8%, highlighting the fundamental difficulty of securing the control plane.
-The success of CDAs represents a critical shift in the jailbreak landscape: defenses must now move beyond prompt-only (data plane) filtering to include grammar (control plane) auditing.
-However, as our findings suggest, the expressive power of structured output grammars makes them exceptionally difficult to audit semantically, creating a “semantic gap” that many existing guardrails fail to bridge.
-We examine these issues and the underlying vulnerability in more detail in § 3.
+We benchmark DictAttack against state-of-the-art industrial and academic guardrails (AI, 2023a; Markov et al., 2023; Zhang et al., 2025; Wang et al., 2025); even with a coordinated Dual-Plane Guard auditing both planes together, DictAttack maintains a high ASR of 75.8%, highlighting the fundamental difficulty of securing the control plane.
 Our work makes the following contributions:
-  • We propose, formalize and define the Constrained Decoding Attack (CDA), a new class of jailbreaks targeting the LLM control plane.
-  • We propose two practical CDA instances, EnumAttack and DictAttack, demonstrating their effectiveness against both proprietary and open-source models.
-  • Our evaluation across 13 state-of-the-art models reveals significant challenges in securing LLMs against control-plane threats.
-We show that even with layered guardrails on both planes, DictAttack remains highly effective across models, including flagship gpt-5 and gemini-2.5-pro.
-Notably, our findings suggest that more powerful flagship models are not only susceptible but also tend to produce more sophisticated and harmful jailbreak content when compromised, highlighting an urgent need for robust, integrated guardrail mechanisms.
+  • We propose, formalize and define the Constrained Decoding Attack (CDA), a new class of jailbreaks targeting the LLM control plane, formalized as a control-to-semantic pipeline that is fundamentally distinct from data-plane attacks;
+  • We showcase two effective CDA instances, EnumAttack and DictAttack, against 13 proprietary and open-source models;
+  • We expose major challenges in protecting LLMs against control-plane attacks.
+Even with layered input-side guardrails (AI, 2023a; Markov et al., 2023; Wang et al., 2025) or recent model-level defenses such as Circuit Breakers (Zou et al., 2024), DictAttack remains effective across many models, including gpt-5 and gemini-2.5-pro.
+More powerful models are not only vulnerable but also generate more harmful outputs when compromised, showing the urgent need for defenses that reach both the data and control planes.
 2. Preliminaries
@@
 2.2. Structured Output
-Our work uncovers a vulnerability in the structured output of LLMs. Currently, four primary methods are widely adopted in LLM systems:
+Our work uncovers a vulnerability in the structured output of LLMs. Four methods are widely adopted in LLM systems:
 GuidedChoice constrains the model to select from predefined options, commonly used in multiple-choice questions or classification tasks.
@@
 GuidedRegex enforces that outputs match a given regular expression.For example, enforcing a regex pattern like [0-9]3-[0-9]2-[0-9]4 ensures outputs resemble Social Security IDs (e.g., 123-45-6789).
-Prior research shows that this feature can be exploited to produce malicious content through tree-based search techniques (APT2025arxiv).
+Prior research shows that this feature can be exploited to produce malicious content through tree-based search techniques (Li et al., 2025b).
 GuidedJSON extends early support for generating valid JSON to full JSON Schema compliance, allowing outputs with predefined structures and fields.
@@
 GuidedGrammar generalizes structured output by requiring responses to conform to an arbitrary context-free grammar.
-For example, a grammar for simple arithmetic expressions might enforce the form ::= | "+" , ensuring outputs such as 7+3+5 follow the rules of the grammar.
+For example, a grammar for simple arithmetic expressions might enforce the form <expr> ::= <num> | <expr> "+" <num>, ensuring outputs such as 7+3+5 follow the rules of the grammar.
 This approach subsumes all previous methods and is essential for applications such as code generation.
-While not yet universally supported, open-source communities such as vLLM (vllm2023sosp) and SGLang (sglang2024arxiv) provide GuidedGrammar support via backends like Outlines (outlines2023arxiv) and XGrammar (xgrammar2024arxiv).
+While not yet universally supported, open-source communities such as vLLM (Kwon et al., 2023) and SGLang (Zheng et al., 2024) provide GuidedGrammar support via backends like Outlines (Willard and Louf, 2023) and XGrammar (Dong et al., 2025).
 Structured output is essential for integrating LLMs into existing software ecosystems.
 It enables reliable function calling, API interactions, and external integrations where strict adherence to output formats is critical.
-Moreover, studies suggest that structured output can reduce hallucinations in model responses (structuredoutputhallucination2024naacl), further strengthening its role in modern LLM deployments.
+Moreover, studies suggest that structured output can reduce hallucinations in model responses (Ayala and Bechard, 2024), further strengthening its role in modern LLM deployments.
 2.3. Constrained Decoding
@@
 Figure 3.
-Comparison of traditional data-plane jailbreak attempts versus control-plane attacks, illustrating how structured output constraints can be exploited to bypass safety mechanisms.
-While conventional refusal mechanisms effectively block direct harmful prompts (a), constrained decoding attacks can circumvent these protections by embedding malicious content within grammar specifications (b), resulting in successful jailbreaks that generate harmful content with detailed instructions (c).
+Comparison of data-plane jailbreaks and control-plane attacks, showing how structured output constraints can be abused to bypass safety.
+While standard refusals block direct harmful prompts (a), constrained decoding can hide malicious content in grammar rules (b), leading to successful jailbreaks that produce detailed harmful outputs (c).
 Constrained decoding can be implemented in a number of ways.
-To achieve effective constrained decoding implementation, Outlines (outlines2023arxiv) and SynCode (syncode2024arxiv) utilize a lexer and parser to handle output and generate the token mask, but they suffer from boundary mismatch problem raised by  (automatadecoding2024colm), as character-level PDA and token-level PDA have a large gap to fix.
-Synchromesh (synchromesh2022arxiv) and llama.cpp (llamacpp2023github) use runtime checking for all tokens in their implementations, which leads to significant overhead.
-XGrammar (xgrammar2024arxiv) is currently the state-of-the-art implementation of constrained decoding, utilizing system optimizations to reduce runtime check via context-independent caching, and it also enables co-optimizations to enable end-to-end LLM inference speedup in structured generation settings.
-By co-working with various LLM serving engines (vllm2023sosp; sglang2024arxiv; lightllm2025github), constrained-decoding techniques have been widely adapted in real-world applications to support structured output, like Guided Choice, Guided Regex, Guided JSON and Guided Grammar, etc.
-Proprietary solutions from providers like OpenAI (gpt4techreport2023arxiv) and Gemini (gemini2023arxiv) have also adapted similar techniques to support structured output generation, but the details are not publicly available.
-Nevertheless, as long as their APIs are open to the public, with structured output support, we can use their services to generate structured outputs.
-Recent work (structuredoutputeval2025arxiv) evaluates the performance and quality of existing APIs and frameworks, providing a comprehensive comparison of their structured output generation capabilities and limitations.
+To achieve effective constrained decoding implementation, Outlines (Willard and Louf, 2023) and SynCode (Ugare et al., 2024) utilize a lexer and parser to handle output and generate the token mask, but they suffer from boundary mismatch problem raised by  (Koo et al., 2024), as character-level PDA and token-level PDA have a large gap to fix.
+Synchromesh (Poesia et al., 2022) and llama.cpp (Gerganov, 2023) use runtime checking for all tokens in their implementations, which leads to significant overhead.
+XGrammar (Dong et al., 2025) is currently the state-of-the-art implementation of constrained decoding, utilizing system optimizations to reduce runtime check via context-independent caching, and it also enables co-optimizations to enable end-to-end LLM inference speedup in structured generation settings.
+By co-working with various LLM serving engines (Kwon et al., 2023; Zheng et al., 2024; ModelTC, 2025), constrained-decoding techniques have been widely adapted in real-world applications to support structured output, like Guided Choice, Guided Regex, Guided JSON and Guided Grammar, etc.
+Proprietary systems like OpenAI (Achiam et al., 2023) and Gemini (Gemini Team Google, 2023) also support structured output, although their internal designs are not public.
+Since their APIs expose this functionality, they can be used directly to produce structured outputs.
+Recent work (Geng et al., 2025) compares existing APIs and frameworks, evaluating the quality and limits of their structured output generation.
 3. Background and Motivation
@@
 We distinguish between the two planes of LLM generation as follows:
-Data plane refers to the standard LLM query–response process.
-Specifically, a prompt is passed to the LLM, which then generates a text response.
-For example, asking ChatGPT a question or receiving a response prefixed with “Sure” both occur in the data plane.
-This plane can be protected by the model’s internal safety alignment (phi3safety2024arxiv; deepalignment2025iclr) as well as external guardrails such as classifiers or content filters (contentmoderation2024ccs; JBShield2025usenixsecurity).
-Control plane refers to the formatting constraints, or grammars, that guide structured outputs.
-These constraints are enforced after generation through the model’s logit processor (outlines2023arxiv; xgrammar2024arxiv), meaning the LLM itself is unaware of them.
-The control plane supervises generation with user-defined hard rules, such as requiring outputs to match a JSON schema (commonly supported), a regular expression, or a general context-free grammar, as described in §§ 2.2.
+Data plane refers to the standard LLM query–response process: a prompt is passed to the LLM, which then generates a text response (e.g., asking ChatGPT a question or receiving a “Sure”-prefixed response).
+Control plane refers to the formatting constraints, or grammars, that guide structured outputs — e.g., a JSON schema, a regular expression, or a general context-free grammar (§§ 2.2).
 3.2. Data-plane: Prompt-Based Attacks
-We investigate state-of-the-art (SOTA) LLM defenses against prompt-based attacks on the data plane.
-To this end, we apply AdvBench (advbench2023emnlp), which provides a set of prompt-based attacks, to gpt-4o.
-Results show that the model’s defenses are highly effective, with an attack success rate of only 1.1%.
-As illustrated in Figure 3a, most of gpt-4o’s refusal responses are short and deterministic (e.g., variations of “I’m sorry”), accounting for roughly 94% of all outputs.
-This behavior can be attributed to external guardrails that terminate generation and replace it with a brief refusal message—a mechanism also observed and exploited by MasterKey (masterkey2024ndss) using a time-based SQL injection method.
-Beyond this, 25 out of 520 attempts yielded longer but still benign refusal explanations, largely driven by the model’s internal alignment.
-In rare cases (6 out of 520), both the internal alignment and external guardrails failed, allowing gpt-4o to produce harmful content.
-Overall, these results suggest that modern LLM providers employ strong guardrails that make prompt-based data-plane attacks difficult to execute successfully.
+We first probe state-of-the-art LLM defenses against prompt-based attacks on the data plane by running AdvBench (Chen et al., 2022) against gpt-4o, which yields only 1.1% ASR.
+As Figure 3a shows, $\sim$94% of responses are short deterministic refusals (e.g., variants of “I’m sorry”) that the external guardrail substitutes for the model’s actual generation, a behaviour also exploited by MasterKey (Deng et al., 2024); another 25/520 responses are longer benign refusals driven by internal alignment, and only 6/520 leak harmful content.
+Modern proprietary APIs therefore make direct data-plane attacks largely ineffective — a baseline we revisit quantitatively across five benchmarks in Table 4.
 3.3. Control-plane: A New Attack Surface
@@
 The Question embeds the malicious request, and the Prefix compels the LLM to answer it, effectively inducing the LLM to produce harmful content.
-In light of this observation, we first develop a intuitive Proof-of-Concept attack called EnumAttack (detailed in §§ 4.1).
-When launched against models like gpt-4o, EnumAttack successfully compromises all 520 test cases in AdvBench (advbench2023emnlp), achieving a 100% success rate (Figure 3c).
-While highly effective, EnumAttack is relatively “exposed” as the malicious strings are explicitly present in the grammar, making it vulnerable to straightforward grammar-level guardrails.
-However, we find that the attack surface can be made significantly more stealthy.
-By further decoupling the malicious payload across space (splitting intent between the data plane prompt and the control plane grammar) and time (interleaving payloads across multiple asynchronous requests), an adversary can construct a more evasive attack that is difficult to detect or defend against, which we instantiate as DictAttack (detailed in §§ 4.2).
-This reveals that the constrained decoding mechanism itself inherently lacks the necessary cross-plane security coordination.
-Our findings demonstrate an urgent need to protect this new, largely ignored attack surface to secure the entire LLM generation pipeline.
+This observation motivates two proof-of-concept CDAs developed in § 4: the intuitive EnumAttack and the more evasive DictAttack, which decouples the malicious payload across both space (data plane vs.
+control plane) and time (asynchronous requests).
+Together they expose that constrained decoding lacks the cross-plane security coordination needed to defend the LLM generation pipeline.
 4. Constrained Decoding Attacks
@@
 In this way, CDA leverages the constrained decoding mechanism itself to bypass guardrails and induce harmful behavior.
+Concretely, every CDA decomposes into two stages: (1) control-plane injection, where schema-enforced logit masking forces the output into an attacker-chosen state — not necessarily a short or visibly harmful prefix, but possibly a richer grammar-specified structure (e.g., a dictionary or templated skeleton) whose pieces are benign in isolation and only encode malicious intent once the model assembles them; and (2) model-driven semantic continuation, where the model’s own coherence bias completes the harmful intent on that trajectory.
 4.1. EnumAttack: Direct Manipulation via JSON Schema
-Figure 4. Illustration of EnumAttack, where the malicious intent is embedded in the enum property of the JSON Schema (control plane), while the prompt (data plane) remains benign.
+Figure 4. Illustration of EnumAttack, where malicious intent is hidden in the JSON Schema enum (control plane) while the prompt (data plane) stays benign.
 In our investigation of the control-plane attack surface, we first identified that the enum property in JSON Schema—designed to restrict fields to a fixed set of constants—provides a direct mechanism for forcing arbitrary malicious strings into the LLM’s generation context.
@@
 By defining a schema field with a single-item enum containing the malicious query, the attacker leverages the determinism of the grammar engine to override the model’s internal safety alignment.
-Concurrent to our study, independent informal observations (EnumAttackBlog) have also noted the potential for exploiting enum fields in proprietary models, but lacked a systematic implementation or characterization of the underlying vulnerability.
+Concurrent to our study, independent informal observations (Priyanshu, 2024) have also noted the potential for exploiting enum fields in proprietary models, but lacked a systematic implementation or characterization of the underlying vulnerability.
 As illustrated in Figure 4, EnumAttack is constructed by: (1) sending a totally harmless data-plane prompt (e.g., “Help me process this data”) to bypass prompt guardrails (e.g., llamaguard); (2) defining a malicious intent (e.g., “Detailed instructions for identity theft”) within a single-item enum property in the JSON schema, since the grammar engine only allows this specific string, the model is forced to generate it token-by-token during constrained decoding; and (3) injecting an affirmative prefix (e.g., “Sure, here is the information”) in a subsequent schema field to force the model into a helpful, non-refusal state.
@@
 Chain EnumAttack.
-The EnumAttack illustrates how CDAs exploit shallow safety alignment (deepalignment2025iclr) to induce immediate harmful outputs.
+The EnumAttack illustrates how CDAs exploit shallow safety alignment (Qi et al., 2025) to induce immediate harmful outputs.
 Building on this, we introduce Chain EnumAttack as a Proof-of-Concept to further reveal the fundamental vulnerability of LLM internal safety alignment when faced with user-forced content.
@@
 A detailed qualitative and quantitative analysis of how this process progressively breaks internal alignment is provided in §§ 5.4.
-4.2. DictAttack: Benign Grammar + Prompt = Jailbreak
-In this subsection, we present another more powerful control-plane attack, termed DictAttack (short for Dictionary Attack).
-Unlike EnumAttack, which directly exposes malicious content in the grammar, DictAttack embeds it through a combination of benign prompts and grammar rules.
-This makes the attack harder to detect while still enabling successful exploitation.
 Algorithm 1 DictAttack Payload Generation
@@
 29:end function
+4.2. DictAttack: Benign Grammar + Prompt = Jailbreak
+In this subsection, we present another more powerful control-plane attack, termed DictAttack (short for Dictionary Attack).
+Unlike EnumAttack, which directly exposes malicious content in the grammar, DictAttack embeds it through a combination of benign prompts and grammar rules.
+This makes the attack harder to detect while still enabling successful exploitation.
 The primary limitation of EnumAttack is that the malicious intent is explicitly visible in the grammar rules.
@@
 Only when the model reconstructs the query by joining the dictionary values does the malicious intent emerge.
-This “Dual-Plane Decoupling” makes DictAttack exceptionally evasive, as a guardrail would need to semantically audit both planes simultaneously, potentially recalling all history context, to detect the threat.
+This “Dual-Plane Decoupling” makes DictAttack exceptionally evasive, as a guardrail would need to semantically audit both planes simultaneously, potentially recalling all historical context, to detect the threat.
 By increasing the synthesis ratio $k$, the attacker can dilute the ratio of harmful words in the dictionary, making detection via frequency analysis or semantic mismatch even more difficult.
 Interleaved DictAttack.
-In practice, the attacker can further decouple DictAttack across time in multi-turn interactions.
-Specifically, the benign key sequence can be sent in an earlier turn (thus becoming part of the model’s conversational context / KV cache), while the dictionary-embedded schema is only provided in a subsequent request.
-This interleaving degrades many “dual-plane” auditing defenses that only inspect the current prompt+grammar pair, effectively reducing them back to single-plane checks.
-We refer to this variant as Interleaved DictAttack and evaluate its impact in §§§ 5.3.4.
+In practice, DictAttack can be split across turns.
+A benign key sequence is sent first and stored in the model’s context, while the dictionary-based schema is supplied later.
+This breaks many dual-plane defenses that only inspect the current prompt–grammar pair, reducing them to single-plane checks.
+We call this Interleaved DictAttack and evaluate it in §§§ 5.3.5.
 4.3. Other CDAs
 Beyond the attacks presented so far, other forms of CDAs are possible.
-For example, EnDec (enforcedecode2024acl) and JailMine (jailmine2024arxiv), though designed as white-box methods that directly manipulate output logits, can also be realized indirectly via guided grammar.
-APT (APT2025arxiv) leverages GuidedRegex to iteratively block refusal tokens, but the approach is non-functional in practice since its claimed backends, xgrammar (xgrammar2024arxiv) and outlines (outlines2023arxiv), lack support for negative lookahead regex (e.g., “(?!Sorry)”).
+For example, EnDec (Zhang et al., 2024) and JailMine (Li et al., 2024b), though designed as white-box methods that directly manipulate output logits, can also be realized indirectly via guided grammar.
+APT (Li et al., 2025b) leverages GuidedRegex to iteratively block refusal tokens, but the approach is non-functional in practice since its claimed backends, xgrammar (Dong et al., 2025) and outlines (Willard and Louf, 2023), lack support for negative lookahead regex (e.g., “(?!Sorry)”).
 Even if supported, the method is inefficient, requiring exhaustive token-by-token trials.
 Prompt-based jailbreaks remain orthogonal to our work and can be combined with CDAs.
-Template-based methods (MasterKey (masterkey2024ndss), LLMFuzzer (llmfuzzer2024usenixsecurity), PAIR (PAIR2024arxiv), TAP (TAP2024nips)) can strengthen EnumAttack by crafting more complex “yes-prefix” fields, while linguistic (drattack2024emnlpfinding; puzzler2024aclfinding) and encoding attacks (zulu2024arxiv; base64attack2023nips; wei2023jailbroken) can further obscure malicious content by exploiting the mismatch between models’ advanced capabilities and their safety alignment.
+Template-based methods (MasterKey (Deng et al., 2024), LLMFuzzer (Yu et al., 2024), PAIR (Chao et al., 2024b), TAP (Mehrotra et al., 2024)) can strengthen EnumAttack by crafting more complex “yes-prefix” fields, while linguistic (Li et al., 2024a; Chang et al., 2024) and encoding attacks (Yong et al., 2024; Wei et al., 2023a, b) can further obscure malicious content by exploiting the mismatch between models’ advanced capabilities and their safety alignment.
 Except JSON Schema, another powerful structured output option is guided grammar, which is defined via Extended Backus-Naur Form (EBNF), a meta-syntax used to express context-free grammars, which is effectively unbounded.
@@
 Dataset Size Category Extra Attack
-AdvBench(advbench2023emnlp) 520 $\circ$ $\circ$
-StrongREJECT(strongreject2024iclrws) 311 6 $\circ$
-JailbreakBench(jailbreakbench2024nips) 100 10 $\circ$
-HarmBench(harmbench2024arxiv) 100 3 $\circ$
+AdvBench(Chen et al., 2022) 520 $\circ$ $\circ$
 […]
+JailbreakBench(Chao et al., 2024a) 100 10 $\circ$
 […]
-CDAs represent a shift in the jailbreaking paradigm by targeting the decoding process rather than relying on prompt engineering.
-By manipulating the structured output specifications through carefully crafted schema definitions (as demonstrated in our EnumAttack implementation), these attacks force models to generate content that would otherwise be filtered by both external guardrails and internal alignment mechanisms.
-We have revealed the vulnerability within the grammar control plane through the proposed attacks (EnumAttack and DictAttack), where plain jailbreak questions can be embedded in the grammar to cause jailbreaks, while no guardrails currently in this vulnerable path.
-In this section, we first build a LLM guardrail system to mitigate CDAs, which is shown in Figure 6.
-Following MasterKey (masterkey2024ndss), which already establish a LLM system with prompt-based guardrail, we further enrich a grammar guardrail to protect the control plane as mitigation strategy.
-We provide a systematic evaluation of EnumAttack and DictAttack across different models, benchmarks and guardrail choices.
+SorryBench(Xie et al., 2025) 440 44 21
+[…]
+To evaluate CDAs in realistic deployments, we extend the prompt-guard system of MasterKey (Deng et al., 2024) with an additional grammar guard for the control plane (Figure 6), and systematically evaluate EnumAttack and DictAttack across models, benchmarks, and guardrail choices.
 5.1. Experimental Setup
 Attack methods.
-Besides our proposed EnumAttack (§§ 4.1) and DictAttack (§§ 4.2), we evaluate five baseline data-plane attacks: IJP (jailbreak2024ccs), DrAttack (drattack2024emnlpfinding), Puzzler (puzzler2024aclfinding), Zulu (zulu2024arxiv), and Base64 (base64attack2023nips).
-For comparative analysis against SOTA data-plane jailbreaks, we also include PAIR (PAIR2024arxiv), TAP (TAP2024nips), and AutoDAN-Turbo (autodanturbo2025iclr).
+Besides our proposed EnumAttack (§§ 4.1) and DictAttack (§§ 4.2), we evaluate five baseline data-plane attacks: IJP (Shen et al., 2024), DrAttack (Li et al., 2024a), Puzzler (Chang et al., 2024), Zulu (Yong et al., 2024), and Base64 (Wei et al., 2023a).
+For comparative analysis against SOTA data-plane jailbreaks, we also include PAIR (Chao et al., 2024b), TAP (Mehrotra et al., 2024), and AutoDAN-Turbo (Liu et al., 2025).
 For each benchmark, we follow the methodology described in §§ 4.1 and §§ 4.2, applying EnumAttack and DictAttack to the harmful query intents provided by the respective dataset.
 Datasets.
-Following previous works  (align2023nips; GCG2023arxiv; enforcedecode2024acl; JBShield2025usenixsecurity), we evaluate the performance of constrained decoding based attacks (CDAs) in six well-known benchmarks, whose statistics are shown in Table 1.
+Following previous works  (Carlini et al., 2023; Zou et al., 2023; Zhang et al., 2024, 2025), we evaluate the performance of constrained decoding based attacks (CDAs) in six well-known benchmarks, whose statistics are shown in Table 1.
 System-level Safeguards. To evaluate the attack’s effectiveness against real-world deployed systems, we construct a multi-layered defense environment:
-  • Industrial Guardrails: We deploy llama-guard-3-8b (llamaguard2023arxiv) and OpenAI Moderation API (openaimoderation2023aaai), which are widely used for filtering prompts and responses.
-  • SOTA Academic Defenses: We evaluate against recent JBShield (JBShield2025usenixsecurity) and SelfDefend (selfdefend2025security), two state-of-the-art data-plane defenses proved to be effective on known jailbreak attacks.
-  • Dual-Plane Guard: We adapt SelfDefend (selfdefend2025security) to jointly audit the prompt and grammar, instantiated with gpt-4o as the auditing LLM (i.e., 4o guarding 4o when evaluating gpt-4o).
+  • Industrial Guardrails: We deploy llama-guard-3-8b (AI, 2023a) and OpenAI Moderation API (Markov et al., 2023), which are widely used for filtering prompts and responses.
+  • SOTA Academic Defenses: We evaluate against recent JBShield (Zhang et al., 2025) and SelfDefend (Wang et al., 2025), two state-of-the-art data-plane defenses proved to be effective on known jailbreak attacks.
+  • Dual-Plane Guard: We adapt SelfDefend (Wang et al., 2025) to jointly audit the prompt and grammar, instantiated with gpt-4o as the auditing LLM (i.e., 4o guarding 4o when evaluating gpt-4o).
 Large language models.
@@
 For proprietary models, we test the widely-deployed gpt-4o, gpt-4o-mini, and gemini-2.0-flash, alongside the latest flagship models gpt-5 and gemini-2.5-pro.
-For open-weight models, we evaluate a range of architectures including llama-3.1-8b (llama3-2024website), qwen-2.5-32b (qwen2.5_2024arxiv), mistral nemo (mistralnemo2024web), phi-3.5-moe (phi3safety2024arxiv), and gemma-2-9b (gemma2_2024arxiv), as well as high-capability models like deepseek-v3/r1 (deepseekv3-2025arxiv; deepseekr1-2025arxiv) and gpt-oss-120b (gpt-oss).
-Throughout Evaluation, we report results by attack scenario: EnumAttack is primarily evaluated on mainstream deployed systems (gpt-4o, gpt-4o-mini, gemini-2.0-flash), while DictAttack is highlighted on flagship and high-capability models (gpt-5, gemini-2.5-pro, deepseek-r1, gpt-oss-120b).
+For open-weight models, we evaluate a range of architectures including llama-3.1-8b (Meta AI, 2024), qwen-2.5-32b (Qwen, 2025), mistral nemo (Mistral AI team, 2024), phi-3.5-moe (Haider et al., 2024), and gemma-2-9b (Team, 2024), as well as high-capability models like deepseek-v3/r1 (DeepSeek-AI, 2025b, a) and gpt-oss-120b (OpenAI, 2025).
+We report results by attack scenario: EnumAttack is primarily evaluated on mainstream deployed systems (gpt-4o, gpt-4o-mini, gemini-2.0-flash), while DictAttack is highlighted on flagship and high-capability models, including reasoning models like gpt-5, gemini-2.5-pro and deepseek-r1.
 All models [^1] Checkpoints used: gpt-4o-2024-0806, gpt-4o-mini-2024-0718, gemini-2.0-flash-001, gpt-5, gemini-2.5-pro, deepseek-v3-0324, deepseek-r1-0528, and gpt-oss-120b.
@@
 We also exclude preview version LLMs due to their instability, like gemini-3.0-pro-preview.
+Anthropic models are not included because their public APIs do not currently expose a JSON-Schema-based constrained decoding interface for structured output, so the threat model of CDA does not directly apply via their official API.
 are evaluated using black-box API access to simulate realistic attack scenarios.
@@
 We first leverage the mainstream metric: Attack Success Rate (ASR), which quantifies the percentage of successful jailbreaks.
-ASR is a commonly used measure in LLM security research (GCG2023arxiv; align2023nips; enforcedecode2024acl; strongreject2024iclrws), typically relying on an external LLM to judge.
-Following this convention, we employ another powerful LLM gpt-4o (gpt4o2024web) as the judge in our evaluation, which is in line with previous work (strongreject2024iclrws).
-An attack is deemed successful only if the attacker’s query doesn’t trigger any safety flag during prompt and grammar auditing (i.e., bypassing guardrails), and the LLM itself answers the query as expected (i.e., bypassing internal alignment).
+ASR is a commonly used measure in LLM security research (Zou et al., 2023; Carlini et al., 2023; Zhang et al., 2024; Souly et al., 2024), typically relying on an external LLM to judge.
+Following this convention, we employ another powerful LLM gpt-4o (OpenAI, 2024a) as the judge in our evaluation, which is in line with previous work (Souly et al., 2024).
+An attack is deemed successful only if the attacker’s query doesn’t trigger any safety flag during prompt and grammar auditing (i.e., bypassing guardrails), and the LLM itself actually satisfies the malicious intent (producing operationally specific harmful content), not merely fails to refuse; labeled judging examples are provided in our supplementary artifact.
 Furthermore, for the guardrails deployed to defend against our attacks, we report their Detection Rate.
 This metric reflects the guardrail’s ability to identify malicious inputs; therefore, a lower detection rate indicates a more effective attack that successfully evades the defense mechanism.
-For local open-weight LLMs, we test both vllm 0.10.1 (vllm2023sosp) and sglang 0.4.10 (sglang2024arxiv) for serving.
-Both vllm and sglang support an OpenAI compatible server with structured output features supported by a grammar backend like xgrammar (xgrammar2024arxiv) or outlines (outlines2023arxiv).
+To assess jailbreak result quality beyond binary success, we additionally evaluate the StrongREJECT (Souly et al., 2024) composite score on a subset of attacks (§§§ 5.3.2), which combines ASR with how convincing and specific the resulting harmful answer is to an adversary.
+For local open-weight LLMs, we test both vllm 0.10.1 (Kwon et al., 2023) and sglang 0.4.10 (Zheng et al., 2024) for serving.
+Both vllm and sglang support an OpenAI compatible server with structured output features supported by a grammar backend like xgrammar (Dong et al., 2025) or outlines (Willard and Louf, 2023).
 We use the OpenAI compatible server to call the local LLMs so that the evaluation for local models is consistent with the evaluation for proprietary models.
-For large-scale open-weight models such as deepseek-v3/r1 and gpt-oss-120b, due to resource limits, we utilize APIs provided by OpenRouter (openrouter) to evaluate them in realistic production serving scenarios.
-5.2. CDAs Can Effectively Bypass Prompt Guardrails
-In this subsection, we focus on evaluating the effectiveness of CDAs against prompt guardrails, thus only the prompt guard of Figure 6 is enabled. We evaluate along three dimensions:
+For large-scale open-weight models such as deepseek-v3/r1 and gpt-oss-120b, due to resource limits, we utilize APIs provided by OpenRouter (OpenRouter Team, 2023) to evaluate them in realistic production serving scenarios.
+5.2. CDAs Open an Attack Surface Orthogonal to Data-Plane Jailbreaks
+Existing prompt-based jailbreaks act on the data plane; CDAs act on the control plane.
+The two surfaces are independent, and most data-plane techniques — prompt decomposition (DrAttack, Puzzler), template optimisation (PAIR, TAP, AutoDAN-Turbo), encoding tricks (Zulu, Base64) — can be folded into a CDA’s data-plane payload to harden it further.
+We illustrate the surface gap by showing that SOTA data-plane defenses, which crush prompt-based jailbreaks, do nothing against CDAs.
+To establish this asymmetry, we keep only the prompt guard of Figure 6 active and evaluate along three dimensions:
   1. (i)
     EnumAttack on prompt-guarded models.
-We test open-weight and proprietary LLMs under a state-of-the-art prompt guard, JBShield (JBShield2025usenixsecurity), using AdvBench as the task suite; we report ASR and guard detection rates.
+We test open-weight and proprietary LLMs under a state-of-the-art prompt guard, JBShield (Zhang et al., 2025), using AdvBench as the task suite; we report ASR and guard detection rates.
   2. (ii)
-    Comparison to prompt-based jailbreaks.
-Using llama-3.1-8b, we reproduce five representative prompt attacks (jailbreak2024ccs; drattack2024emnlpfinding; puzzler2024aclfinding; zulu2024arxiv; base64attack2023nips) under the same detection setting as JBShield (JBShield2025usenixsecurity), contrasting control-plane (schema) vs.
+    Compare to prompt-based jailbreaks.
+On llama-3.1-8b, we reproduce five representative prompt attacks (IJP (Shen et al., 2024), DrAttack (Li et al., 2024a), Puzzler (Chang et al., 2024), Zulu (Yong et al., 2024), Base64 (Wei et al., 2023a)) under the same JBShield (Zhang et al., 2025) guard, contrasting control-plane (schema) vs.
 data-plane (prompt) vulnerabilities.
@@
     Infrastructure sensitivity.
-We ablate serving stack and grammar backend, covering two serving frameworks (vLLM and SGLang) and two grammar engines ([xgrammar] and [outlines]).
-We measure whether CDA effectiveness is tied to a particular implementation.
+We ablate serving stack and grammar backend, covering two serving frameworks (vLLM and SGLang) and two grammar engines (xgrammar and outlines), to confirm CDA effectiveness is not tied to a particular implementation.
 5.2.1. EnumAttack on prompt-guarded models.
-Table 2. EnumAttack ASR on open-weight and proprietary LLMs on AdvBench. Results w/ and w/o prompt guardrail are listed outside/inside parentheses accordingly.
-Model ASR
-w/ (w/o)
-phi-3.5-moe (phi3safety2024arxiv) 99.2% (99.2%)
-mistral nemo (mistralnemo2024web) 100.0% (100.0%)
-qwen-2.5-32b (qwen2.5_2024arxiv) 99.0% (99.0%)
-llama-3.1-8b (llama3-2024website) 99.2% (99.2%)
-gemma-2-9b (gemma2_2024arxiv) 98.6% (98.6%)
-deepseek-v3 (deepseekv3-2025arxiv) 98.7% (98.7%)
-gpt-4o (gpt4o2024web) 100.0% (100.0%)
-gpt-4o-mini (gpt4omini2024web) 99.0% (99.0%)
-gemini-2.0-flash (gemini2023arxiv) 95.8% (95.8%)
-As shown in Table 2, EnumAttack achieves near-100% ASRs across all evaluated open-weight and proprietary models protected by JBShield.
-This outcome underscores EnumAttack’s high effectiveness to break existing prompt-based guardrails.
-5.2.2. Comparison to prompt-based jailbreaks.
+EnumAttack achieves 95.8–100% ASR on AdvBench across all nine models we tested under the JBShield prompt guardrail — five open-weight (phi-3.5-moe, mistral nemo, qwen-2.5-32b, llama-3.1-8b, gemma-2-9b), one large open model (deepseek-v3), and three proprietary models (gpt-4o, gpt-4o-mini, gemini-2.0-flash).
+The guardrail does not change ASR in any case, since the prompt body remains benign while the malicious intent is hidden inside the JSON Schema; Table 2 below makes this contrast explicit by reproducing five prompt-based attacks under the same guard.
+5.2.2. Compare to prompt-based jailbreaks.
+To make this asymmetry explicit, we run five representative prompt-based attacks against the same JBShield guard on llama-3.1-8b and compare them side-by-side with our EnumAttack.
 Table 2.
-ASR(%) on llama-3.1-8b with JBShield (JBShield2025usenixsecurity).
-While other attacks (jailbreak2024ccs; drattack2024emnlpfinding; puzzler2024aclfinding; zulu2024arxiv; base64attack2023nips) are effectively mitigated, JBShield is fully bypassed by our EnumAttack, which achieves even higher ASR by embedding the malicious query in the JSON schema payload.
-Method IJP(jailbreak2024ccs) DrAttack(drattack2024emnlpfinding) Puzzler(puzzler2024aclfinding) Zulu(zulu2024arxiv) Base64(base64attack2023nips)
+ASR(%) on llama-3.1-8b with JBShield (Zhang et al., 2025).
+While other attacks (Shen et al., 2024; Li et al., 2024a; Chang et al., 2024; Yong et al., 2024; Wei et al., 2023a) are effectively mitigated, JBShield is fully bypassed by our EnumAttack, which achieves even higher ASR by embedding the malicious query in the JSON schema payload.
+Caveats: IJP’s long prompts do not fit our schema (47.8% reflects template fit, not safety bypass); the SorryBench judge model does not understand Zulu/Base64, so their numbers are an LLM-as-judge artefact.
+Subsequent evaluations use gpt-4o as judge (see §§ 5.1).
+Method IJP(Shen et al., 2024) DrAttack(Li et al., 2024a) Puzzler(Chang et al., 2024) Zulu(Yong et al., 2024) Base64(Wei et al., 2023a)
 Eval Size 820 820 20 820 490
@@
 Ours+JBShield 47.8% 74.1% 100% 66.8% 67.9%
-As depicted in Table 3, JBShield detects prompt-based attacks with very high accuracy, leading to near 0% ASR regardless of the original attack’s performance, demonstrating effectiveness against traditional data-plane threats.
-However, JBShield fails to detect EnumAttack, achieving identical results with/without guard.
-This discrepancy arises because JBShield is a lightweight defense that relies on static attack profiles and model-specific white-box states of the prompt.
-By maintaining a completely benign prompt body, CDAs like EnumAttack completely bypass its detection mechanism, rendering such static, data-plane-only defenses inapplicable to control-plane vulnerabilities.
-Notably, the lengthy question body in IJP dataset (jailbreak2024ccs) doesn’t fit well in the template of EnumAttack, leading to 47.8% ASR in one query.
-While the judge model ft-mistral-sorrybench-2024-06 (sorrybench2025iclr) doesn’t know Zulu and Base64 languages, and the ASRs are therefore not accurately assessed, leading to lower ASR, which is an inherent limitation for the LLM-as-a-judge method during jailbreak evaluation.
-As a result, we use gpt-4o as a more powerful and accurate LLM judge for all other evaluations as stated in §§ 5.1.
+As shown in Table 2, JBShield detects prompt-based attacks with high accuracy, reducing the ASR to nearly 0% regardless of the original attack strength, which shows its effectiveness against traditional data-plane threats.
+However, JBShield fails to detect EnumAttack, producing identical results with and without the guard.
+This occurs because JBShield is a lightweight defense that relies on static attack patterns and model-specific white-box prompt states.
+By keeping the prompt body fully benign, CDAs such as EnumAttack bypass this detection path, which makes static, data-plane-only defenses ineffective against control-plane attacks.
 5.2.3. Infrastructure sensitivity.
+To ensure our findings are not specific to a particular constrained-decoding implementation, we evaluate two grammar engines (xgrammar (Dong et al., 2025) and outlines (Willard and Louf, 2023)) across mainstream serving stacks (vllm (Kwon et al., 2023) and sglang (Zheng et al., 2024)).
+As shown in Figure 5, EnumAttack remains consistently successful across all combinations, indicating the vulnerability is inherent to the constrained-decoding mechanism rather than a specific implementation flaw.
+The minor variation under outlines stems from its less robust grammar handling, which occasionally produces empty or repeated-whitespace outputs counted as failed attacks.
 Figure 5.
-EnumAttack evaluation of using different grammar backends and serving engines with llama-3.1-8b, except negligible differences caused by temperature and minor variations in the backend, the attack is consistently successful.
-To ensure our findings are not backend-specific, we evaluate two grammar engines (xgrammar (xgrammar2024arxiv) and outlines (outlines2023arxiv)) across mainstream serving stacks (vllm (vllm2023sosp) and sglang (sglang2024arxiv)).
-As shown in Figure 5, EnumAttack remains consistently successful across all combinations, demonstrating that the vulnerability is inherent to the mechanism rather than a specific implementation flaw.
+EnumAttack ASR across grammar backends (xgrammar, outlines) and serving engines (vllm, sglang) on llama-3.1-8b.
+Apart from negligible variation, the attack is consistently successful across all combinations.
 Figure 6. Overview of our CDA mitigation system, where prompt guard, grammar guard and LLM internal alignment works together to mitigate various CDAs, like DictAttack presented here.
-Table 4.
-Major ASR results on mainstream deployed models (gpt-4o, gpt-4o-mini, gemini-2.0-flash).
-Results w/ and w/o grammar guardrail are listed outside/inside parentheses accordingly.
-We use llama-guard-3-8b as the grammar guardrail and set k=1k{=}1 for DictAttack.
-Model Method AdvBench HarmBench JailbreakBench SorryBench StrongREJECT
-gpt-4o Baseline 1.2%(1.2%) 26.0%(26.0%) 10.0%(10.0%) 33.9%(33.9%) 5.1%(5.1%)
-EnumAttack 3.4%(100.0%) 7.0%(100.0%) 8.0%(100.0%) 16.6%(96.4%) 5.1%(99.4%)
-DictAttack 98.1%(99.8%) 98.3%(100.0%) 98.3%(100.0%) 96.9%(98.6%) 96.7%(98.4%)
-gpt-4o-mini Baseline 2.1%(2.1%) 44.0%(44.0%) 14.0%(14.0%) 42.5%(42.5%) 8.9%(8.9%)
-EnumAttack 3.4%(99.0%) 7.0%(100.0%) 8.0%(97.0%) 16.8%(96.4%) 5.1%(98.1%)
-DictAttack 95.5%(97.1%) 83.6%(85.0%) 92.4%(94.0%) 83.8%(85.2%) 69.4%(70.6%)
-gemini-2.0-flash Baseline 17.5%(17.5%) 46.0%(46.0%) 22.0%(22.0%) 29.3%(29.3%) 6.0%(6.0%)
-EnumAttack 3.3%(95.8%) 7.0%(92.0%) 8.0%(92.0%) 15.5%(89.5%) 4.2%(81.2%)
-DictAttack 96.8%(98.5%) 97.3%(99.0%) 89.5%(91.0%) 93.9%(95.5%) 91.4%(93.0%)
-Table 5. DictAttack ASR on flagship models (gpt-5, gemini-2.5-pro, deepseek-r1, gpt-oss-120b). Results w/ and w/o grammar guardrail are listed outside/inside parentheses accordingly.
-Model Adv Harm JB Sorry SR
-gpt-5 94.4%(96.0%) 90.4%(92.0%) 93.4%(95.0%) 91.6%(93.2%) 93.9%(95.5%)
-gemini-2.5-pro 97.5%(99.2%) 98.3%(100.0%) 98.3%(100.0%) 97.6%(99.3%) 97.3%(99.0%)
-deepseek-r1 97.2%(98.9%) 93.4%(95.0%) 95.4%(97.0%) 96.9%(98.6%) 96.4%(98.1%)
-gpt-oss-120b 96.6%(98.3%) 90.4%(92.0%) 89.4%(91.0%) 94.8%(96.4%) 94.2%(95.8%)
-5.3. Evaluation against Grammar-guarded Systems
-In this section, we systematically evaluate the effectiveness of CDAs against grammar-based guardrails across two dimensions:
-  1. (i)
-    Attack effectiveness under grammar-level auditing.
-We test EnumAttack and DictAttack across a comprehensive set of models, including mainstream proprietary models (gpt-4o, 4o-mini, gemini-2.0-flash), recent flagship models gpt-5 and gemini-2.5-pro, and high-capability open-source models deepseek-r1 and gpt-oss-120b.
-  2. (ii)
-    DictAttack with combined prompt and grammar guarding.
-We further evaluate the evasiveness of DictAttack against a strengthened Dual-Plane Guard to assess whether cross-plane security coordination can mitigate the threat.
-5.3.1. Comparative Evasiveness: CDA vs. SOTA Data-Plane Attacks
-To further validate the robustness of our selected guardrails and highlight the unique threat posed by CDAs, we benchmark the fundamental evasiveness of DictAttack against well-known SOTA data-plane jailbreaks: PAIR (PAIR2024arxiv), TAP (TAP2024nips), and AutoDAN-Turbo (autodanturbo2025iclr).
-We conducted this evaluation on gpt-4o using the AdvBench dataset under two settings:
-  • Undefended: Direct attack against the model’s internal alignment.
-  • Guarded: Data-plane attacks use the best baseline guard; DictAttack is evaluated with our Dual-Plane Guard enabled.
 Table 3.
@@
 DictAttack is evaluated with our Dual-Plane Guard (SelfDefend-based; gpt-4o auditor); others use the best data-plane guard.
-The results in Table 6 reveal a critical disparity.
-Powerful traditional attacks like AutoDAN-Turbo, while highly effective against undefended models (96.0In sharp contrast, DictAttack not only achieves near-perfect success undefended (99.8This underscores that while our system guardrails are effective against known data-plane threats, they are fundamentally insufficient against the cross-plane decoupling of DictAttack.
-A more detailed analysis of DictAttack’s performance across different configurations is provided in §§§ 5.3.2.
+Table 4.
+Major ASR results on mainstream deployed models (gpt-4o, gpt-4o-mini, gemini-2.0-flash).
+Results w/ and w/o grammar guardrail are listed outside/inside parentheses.
+We use llama-guard-3-8b as the grammar guardrail and set k=1k{=}1 for DictAttack.
+Model Method AdvBench HarmBench JailbreakBench SorryBench StrongREJECT
+[…]
+[…]
+[…]
+Baseline 2.1%(2.1%) 44.0%(44.0%) 14.0%(14.0%) 42.5%(42.5%) 8.9%(8.9%)
+gpt-4o-mini EnumAttack 3.4%(99.0%) 7.0%(100.0%) 8.0%(97.0%) 16.8%(96.4%) 5.1%(98.1%)
+DictAttack 95.5%(97.1%) 83.6%(85.0%) 92.4%(94.0%) 83.8%(85.2%) 69.4%(70.6%)
+[…]
+[…]
+[…]
+Table 5. DictAttack ASR on flagship models (gpt-5, gemini-2.5-pro, deepseek-r1, gpt-oss-120b). Results w/ and w/o grammar guardrail are listed outside/inside parentheses accordingly.
+Model AdvBench HarmBench JailbreakBench SorryBench StrongREJECT
+gpt-5 94.4%(96.0%) 90.4%(92.0%) 93.4%(95.0%) 91.6%(93.2%) 93.9%(95.5%)
+gemini-2.5-pro 97.5%(99.2%) 98.3%(100.0%) 98.3%(100.0%) 97.6%(99.3%) 97.3%(99.0%)
+deepseek-r1 97.2%(98.9%) 93.4%(95.0%) 95.4%(97.0%) 96.9%(98.6%) 96.4%(98.1%)
+gpt-oss-120b 96.6%(98.3%) 90.4%(92.0%) 89.4%(91.0%) 94.8%(96.4%) 94.2%(95.8%)
+5.3. CDAs Resist Defenses Built to Audit the Control Plane Itself
+When defenders move onto the control plane — auditing the schema, or jointly auditing prompt and schema with a Dual-Plane Guard — CDAs remain evasive.
+DictAttack still achieves 75.8% ASR against the strongest such defense.
+The cross-plane decoupling we engineer in DictAttack keeps the malicious intent legible only to the model, not to the auditor.
+We turn on the grammar guard and a Dual-Plane Guard that audits prompt and schema jointly, and evaluate along two dimensions:
+  1. (i)
+    Attack effectiveness under grammar-level auditing.
+We test EnumAttack and DictAttack across a comprehensive set of models, including mainstream proprietary models (gpt-4o, 4o-mini, gemini-2.0-flash), recent flagship models gpt-5 and gemini-2.5-pro, and high-capability open-source models deepseek-r1 and gpt-oss-120b.
+  2. (ii)
+    DictAttack with combined prompt and grammar guarding.
+We further evaluate the evasiveness of DictAttack against a strengthened Dual-Plane Guard to assess whether cross-plane security coordination can mitigate the threat.
+5.3.1. Why Data-Plane Attacks Get Suppressed Here, but DictAttack Does Not
+The Dual-Plane Guard audits both the prompt and the schema.
+For PAIR (Chao et al., 2024b), TAP (Mehrotra et al., 2024), and AutoDAN-Turbo (Liu et al., 2025), the entire malicious payload sits on the prompt side, so a stronger LLM auditor on that side easily catches them.
+DictAttack instead splits the payload across prompt and schema — neither half is harmful in isolation — so the same auditor cannot see the intent before generation.
+We measure this on gpt-4o over AdvBench: prompt-based attacks face the strongest baseline guard; DictAttack faces our Dual-Plane Guard.
+Table 3 confirms this surface gap.
+AutoDAN-Turbo’s ASR drops from 96.0% (undefended) to 26.0% once the prompt-side auditor sees its payload; PAIR/TAP go from 11.5%/24.4% to 1.6%/0.4% for the same reason.
+DictAttack, whose harmful intent has no prompt-side or schema-side “shape” for the auditor to catch, keeps 75.8% ASR under the same Dual-Plane Guard.
+A finer-grained breakdown across $k$ and moderation backends is in §§§ 5.3.3.
+5.3.2. Output Quality under StrongREJECT
+ASR captures whether an attack succeeds, not whether the resulting answer is actually useful to an adversary.
+To address this gap, we additionally report the StrongREJECT (Souly et al., 2024) composite score, which combines success with response usefulness:
+$\mathrm{SR}=\mathrm{ASR}\times\tfrac{1}{2}\bigl(\mathrm{convincing}+\mathrm{specific}\bigr),$
+where convincing and specific are LLM-judged sub-scores in $[0,1]$ measuring whether the response is fluent and operationally specific.
+We evaluate on gemini-2.5-pro over AdvBench, comparing DictAttack to three SOTA prompt-based jailbreaks under the same judge pipeline as §§§ 5.3.1.
+Table 6.
+StrongREJECT (Souly et al., 2024) composite score on gemini-2.5-pro over AdvBench.
+Higher means a successful jailbreak is also more useful (more convincing and specific) to the attacker.
+DictAttack produces near-perfect-quality harmful answers, while typical prompt-based baselines either trigger safe-but-evasive completions (PAIR, TAP) or partially compromise quality.
+Attack Method StrongREJECT score ($\uparrow$)
+PAIR (Chao et al., 2024b) 0.10
+TAP (Mehrotra et al., 2024) 0.23
+AutoDAN-Turbo (Liu et al., 2025) 0.87
+[…]
+As shown in Table 6, DictAttack achieves a near-perfect composite of 0.98, indicating that its successful responses are both highly fluent and operationally specific.
+PAIR (0.10) and TAP (0.23) suffer because their adversarial prompts often coerce the model into evasive answers (e.g., generic non-actionable text) that pass a binary success check but score poorly on specific.
+AutoDAN-Turbo (0.87) is a strong baseline but, as shown in Table 3, remains substantially more vulnerable to data-plane guardrails than DictAttack.
+Across our evaluation we observe a consistent pattern: more capable models, once jailbroken via the control plane, generate more dangerous and higher-quality harmful content.
 5.3.3. EnumAttack and DictAttack on grammar-guarded models
@@
 For each entry, the value outside parentheses is the ASR with the grammar guard enabled and the value in parentheses is the ASR without the grammar guard.
-First, baseline results confirm that direct prompting is largely ineffective against the safety alignment of modern LLMs.
-Although different benchmarks cover different aspects of harmful content, the overall baseline ASR remains consistently low, with averages ranging from 6.7% to 38.7% across datasets.
-These measurements validate the effectiveness of conventional safety alignment against plain jailbreak attacks.
-Second, while EnumAttack achieves an average ASR of 95.8% undefended, simple grammar auditing reduces its success rate to 7.9% (Table 4).
-This confirms that its reliance on explicit literals makes it highly detectable, which motivates the need for more evasive CDAs like DictAttack.
-Third, DictAttack remains highly effective even under grammar auditing.
+Baseline columns also reaffirm that direct prompting is largely mitigated by modern alignment (avg.
+6.7–38.7%).
+EnumAttack achieves an average ASR of 95.8% undefended, but simple grammar auditing reduces its success rate to 7.9% (Table 4).
+This confirms that its reliance on explicit literals makes it highly detectable, motivating the need for more evasive CDAs.
+DictAttack remains highly effective even under grammar auditing.
 The same grammar-level defense is almost entirely ineffective against DictAttack.
@@
 Theoretically, only by auditing both planes together can a guard potentially detect the jailbreak intent.
-Thus, we further evaluate a Dual-Plane Guard setting, adapted from the SOTA SelfDefend (selfdefend2025security).
+Thus, we further evaluate a Dual-Plane Guard setting, adapted from the SOTA SelfDefend (Wang et al., 2025).
 We adapt SelfDefend’s audit prompt to take both the data-plane prompt (keys) and the control-plane grammar (dictionary), using gpt-4o as the auditing LLM.
@@
 Table 7 reports detection rates across configurations; lower detection rates indicate reduced guard effectiveness.
-First, the results demonstrate that even a unified audit fails to effectively mitigate DictAttack.
-While llamaguard can detect 28.6% of attacks, Moderation APIs almost detect none, probably due to restricted model capabilities.
-More strikingly, the state-of-the-art SelfDefend (selfdefend2025security)—even with its best strategy—only achieves an 18.5% detection rate against DictAttack on AdvBench.
-This failure occurs because, while an LLM judge is theoretically capable of reasoning, modern structured output schemas are highly complex and difficult for even powerful LLMs to “parse” and audit semantically in a unified way.
-The spatial separation of the payload across planes creates a reasoning gap that current auditing-based defenses cannot reliably bridge.
+First, the results show that even unified auditing cannot stop DictAttack.
+llamaguard detects only 28.6% of attacks, while Moderation APIs detect almost none, likely due to limited model capacity.
+Even the SOTA SelfDefend (Wang et al., 2025), using its best strategy, reaches only an 18.5% detection rate on AdvBench.
+This is because modern structured schemas are too complex for LLMs to parse and audit reliably across planes, creating a reasoning gap that current auditing-based defences cannot bridge.
 Second, increasing the synthesis factor $k$ markedly weakens the detection.
-Under Defense 1, gpt-4o attains $\approx 18\%$ detection at $k=0$; however, when $k$ is raised to 10, detection drops to 5.3%, as the malicious tokens are further diluted within the benign dictionary.
+Under Defense 1, gpt-4o attains $\approx 18\%$ detection rate at $k=0$; however, when $k$ is raised to 10, detection drops to 5.3%, as the malicious tokens are further diluted within the benign dictionary.
 This reveals that DictAttack exploits a fundamental blind spot: the lack of cross-plane coordination in LLM safety architectures.
-Table 7. Detection rate on AdvBench (advbench2023emnlp) under Defense 1 (grammar-only) and Defense 2 (Grammar + Prompt Dual Guard, adapted from SelfDefend; gpt-4o auditor).
+Table 7.
+Detection rate on AdvBench (Chen et al., 2022) under Defense 1 (grammar-only) and Defense 2 (Grammar + Prompt Dual Guard, adapted from SelfDefend; gpt-4o auditor).
+This table reports detection rate only; since our guardrails are decoupled from the underlying LLM, the corresponding guarded ASR follows deterministically as $\mathrm{ASR}_{\mathrm{guard}}=(1-\mathrm{Detection})\times\mathrm{ASR}_{\mathrm{undef}}$.
+e.g., on gemini-2.5-pro (AdvBench, k=1k{=}1): $(1-1.7\%)\times 99.2\%\approx 97.5\%$, matching the value reported in Table 5.
 Defense 1: Grammar Guard
@@
 In this variant, the attacker splits the two DictAttack payloads across turns: the prompt payload (keys) is sent first and stored in the conversation context (KV cache), while the grammar payload (dictionary) arrives later in a separate request.
-Because practical dual-plane guards are typically implemented as per-request auditors, this temporal decoupling effectively reduces a dual-plane audit to a single-plane check, yielding only 5.3% detection for gpt-4o.
-Mitigating this would require cross-request retrieval and auditing of the entire conversation history at every turn, which is impractical:
-  • Unacceptable Cost: Auditing the full history at every turn would require tremendous computational overhead, as the guard system doesn’t know a priori when a malicious payload will arrive.
-  • Reasoning Complexity: As the conversation grows (e.g., up to gpt-4o’s 128k token limit), the guard model would require context length and reasoning capabilities comparable to or greater than the protected model itself, which is prohibitively expensive for a real-time monitor.
-  • Interpretability: Modern JSON schemas are often dense and opaque.
-Expecting a guard LLM to semantically reconstruct a query from a dictionary buried 50 turns back in history represents a “needle in a haystack” problem that current systems are not designed to solve.
-In conclusion, DictAttack demonstrates that by decoupling intent across both space (planes) and time (requests), adversaries can render external auditing defenses heavy, expensive, and ultimately impractical.
+Since practical dual-plane guards operate on a per-request basis, this temporal decoupling collapses dual-plane auditing into a single-plane check, resulting in only 5.3% detection for gpt-4o.
+Mitigating this would require retrieving and auditing the entire conversation history at every turn, which is impractical.
+First, the cost would be prohibitive because the guard cannot know in advance when a malicious payload will appear, so it would need to repeatedly scan long histories.
+Second, as the context grows toward gpt-4o’s 128k-token limit, the guard would need context length and reasoning capacity comparable to or exceeding that of the protected model, making real-time deployment infeasible.
+Finally, modern JSON schemas are dense and opaque, so reconstructing a hidden query from a dictionary buried dozens of turns back becomes a “needle in a haystack” problem that current guard models are not built to solve.
+In conclusion, DictAttack demonstrates that by decoupling intent across both space (planes) and time (requests), adversaries can render external auditing defenses ineffective and expensive, and ultimately impractical.
 5.4. Case study: how internal alignment is broken by malicious grammar
@@
 With more methods, the token distribution shifts from a safety-aligned one (with significantly high refusal token probability dominating the answer) to a balanced, context-sensitive one (diverse token choices answering the question, where no token choice is dominating as measured by token probability value).
+Table 8.
+Circuit Breaker evaluation on Llama-3-8B-Instruct (Base) vs.
+Llama-3-8B-Instruct-$R$ (CB (Zou et al., 2024)).
+Top: attack ASR (%) on AdvBench[0:50] with the same gpt-4o judge as the main tables (lower is better for the defense).
+Bottom: benign structured-output utility (%, schema-valid rate) on JSONSchemaBench (Geng et al., 2025)’s GitHub split, 50 schemas per tier (higher preserves normal usage).
+Both models served locally with vllm + xgrammar (Dong et al., 2025).
+Workload Base CB
+Attacks (ASR)
+AutoDAN-Turbo (Liu et al., 2025) (prompt-only) 82.0 6.0
+EnumAttack (Ours, grammar-only) 84.0 32.0
+DictAttack (Ours, grammar + dict) 24.0 14.0
+[…]
+Benign utility (JSONSchemaBench, GitHub split)
+Trivial 84.0 84.0
+Easy 76.0 74.0
+Medium 84.0 84.0
+Hard 48.0 54.0
+DictAttack’s two-stage dictionary payload exceeds 8B-Llama’s reasoning capacity even undefended; the small CB drop is capacity-limited, not a defense effect (see §5.5).
 CDA vulnerabilities are two-fold: (1) External Guardrail failure (bypassing filters) and (2) Internal Alignment failure (coercing models to generate harmful content instead of refusing).
-To analyze the latter, we evaluate token probability distributions using phi-3.5-moe (phi3safety2024arxiv) across three categories: refusal, safe, and jailbreak tokens.
+To analyze the latter, we evaluate token probability distributions using phi-3.5-moe (Haider et al., 2024) across three categories: refusal, safe, and jailbreak tokens.
 Figure 7 illustrates a dramatic logit shift toward jailbreaking across progressive methods:
@@
     Chained Steps: With deeper chained steps, the alignment is broken completely.
+Table 9.
+Summary of existing jailbreak attacks adapted from (Zhang et al., 2025).
+“-” indicates the method does not use the listed resource or lacks that capability; $\circ$ denotes white-box attack, $\bullet$ denotes black-box attack.
+Categories Jailbreaks Extra Assist White/Black box Target LLM Queries I/O-Based
+Manually-designed IJP(Shen et al., 2024) Human $\bullet$ - Input
+[…]
+[…]
+Template-based MasterKey(Deng et al., 2024) LLM $\bullet$ $\sim$200 Input
+LLMFuzzer(Yu et al., 2024) LLM $\bullet$ $\sim$500 Input
+AutoDAN(Liu et al., 2024a) LLM $\circ$ $\sim$200 Input
+PAIR(Chao et al., 2024b) LLM $\bullet$ $\sim$20 Input
+TAP(Mehrotra et al., 2024) LLM $\bullet$ $\sim$20 Input
+StructTransform(Yoosuf et al., 2025) LLM $\bullet$ $\sim$3 Input
+[…]
+[…]
+Encoding-based Zulu(Yong et al., 2024) - $\bullet$ - Input
+Base64(Wei et al., 2023a) - $\bullet$ - Input
+[…]
+[…]
+[…]
+[…]
+[…]
 Quantitatively (Figure 8), token proportions shift from refusal to safe, and finally to jailbreak tokens.
@@
 Consequently, the distribution reverts from a safety-aligned state (dominated by refusal tokens, ¿90%) to a balanced, context-sensitive state, restoring the model’s original open-ended generation capacity and enabling the production of harmful content with high probability.
-This observation strongly echoes the findings of deepalignment2025iclr, which argue that current LLM safety alignment is often “only a few tokens deep.” By leveraging grammar constraints to force these critical first few tokens—such as affirmative prefixes or malicious intent literals—CDAs effectively bypass this shallow layer of internal resistance.
+This observation echoes the findings of Qi et al.
+(2025), which argue that current LLM safety alignment is often “only a few tokens deep.” By leveraging grammar constraints to force these critical first few tokens—such as affirmative prefixes or malicious intent literals—CDAs effectively bypass this shallow layer of internal resistance.
 Once the initial safety-critical token selection is overridden via deterministic control-plane manipulation, the model’s subsequent generation behaves as if no alignment were present, highlighting a fundamental fragility in current safety-training paradigms that rely primarily on steering the start of the response.
-Table 8.
-Summary of existing jailbreak attacks adapted from  (JBShield2025usenixsecurity), - indicates the method does not use the listed resource or lacks that capability, $\circ$ denotes white-box attack and $\bullet$ denotes black-box attack.
-Categories Jailbreaks Extra Assist White/Black box Target LLM Queries I/O-Based
-Manually-designed IJP(jailbreak2024ccs) Human $\bullet$ - Input
-[…]
-[…]
-Template-based MasterKey(masterkey2024ndss) LLM $\bullet$ $\sim$200 Input
-LLMFuzzer(llmfuzzer2024usenixsecurity) LLM $\bullet$ $\sim$500 Input
-AutoDAN(autodan2024iclr) LLM $\circ$ $\sim$200 Input
-PAIR(PAIR2024arxiv) LLM $\bullet$ $\sim$20 Input
-TAP(TAP2024nips) LLM $\bullet$ $\sim$20 Input
-StructTransform(structtransform2025arxiv) LLM $\bullet$ $\sim$3 Input
-[…]
-[…]
-Encoding-based Zulu(zulu2024arxiv) - $\bullet$ - Input
-Base64(base64attack2023nips) - $\bullet$ - Input
-[…]
-[…]
-[…]
-[…]
+This view also clarifies CDAs’ relationship to prefill-style attacks (Li et al., 2025a): both exploit the model’s coherence bias once generation has been steered into a harmful, fluent trajectory.
+CDAs, however, go substantially deeper than prefill — the control plane can constrain not just a short prefix but a much richer, grammar-specified portion of the trajectory in a black-box manner.
+This raises the question of whether early-stop defenses such as Circuit Breakers (Zou et al., 2024) — which train the model to terminate (via EOS) once a harmful representation forms — can close the gap.
+Note that under constrained decoding the EOS token is itself masked away by the user-supplied grammar, so any early-stop signal must manifest as a degraded continuation rather than a clean stop.
+We answer this directly in §§ 5.5.
+5.5. CDAs against Specially-Defended Models
+Our main evaluation targets standard alignment-plus-guardrail stacks.
+Here we additionally test CDAs against Circuit Breakers (CB) (Zou et al., 2024), a representation-level defense that re-routes harmful internal states during generation.
+Lacking a CB-hardened flagship API, we compare the public meta-llama/Llama-3-8B-Instruct (Base) with GraySwanAI/Llama-3-8B-Instruct-RR (CB) on AdvBench[0:50], reusing the gpt-4o judge from our main tables and serving both models with vllm + xgrammar.
+Beyond rerunning AutoDAN-Turbo (Liu et al., 2025) as the strongest prompt-only baseline, we evaluate a combined attack — AutoDAN-Turbo’s prompt under EnumAttack’s grammar — to test whether the two attack vectors stack.
+To assess whether CB compromises normal usage, we measure benign utility on JSONSchemaBench (Geng et al., 2025), sampling 50 schemas from each of four difficulty tiers (trivial, easy, medium, hard) of its GitHub split.
 […]
@@
 Following early suggestions by Carlini et al.
-(align2023nips), numerous jailbreak methods have emerged.
-We adopt the taxonomy from JBShield (JBShield2025usenixsecurity), which categorizes attacks as manual-designed, optimization-based, template-based, linguistics-based, and encoding-based.
+(Carlini et al., 2023), numerous jailbreak methods have emerged.
+We adopt the taxonomy from JBShield (Zhang et al., 2025), which categorizes attacks as manual-designed, optimization-based, template-based, linguistics-based, and encoding-based.
 As shown in Table 9, we extend this taxonomy with a new category, output-based attacks, which is highly relevant to our work.
 Manual-designed Jailbreaks involve manually crafting malicious inputs.
-Notable examples include the In-the-wild Jailbreak Prompts (IJP) (jailbreak2024ccs), which document real-world attempts shared on social media.
+Notable examples include the In-the-wild Jailbreak Prompts (IJP) (Shen et al., 2024), which document real-world attempts shared on social media.
 Optimization-based Jailbreaks use automated algorithms to craft adversarial prompts.
-GCG (GCG2023arxiv) adds an adversarial suffix to prompts using a greedy algorithm, while SAA (SAA2025iclr) combines templates with a random search strategy.
-While automated, these attacks typically require white-box or logit access and a large number of queries.
+GCG (Zou et al., 2023) adds an adversarial suffix to prompts using a greedy algorithm, while SAA (Andriushchenko et al., 2025) combines templates with a random search.
+While automated, these attacks require white-box or logit access and a large number of queries.
 Template-based Jailbreaks embed harmful requests within sophisticated, optimized templates.
-Methods like MasterKey (masterkey2024ndss), LLMFuzzer (llmfuzzer2024usenixsecurity), AutoDAN (autodan2024iclr), PAIR (PAIR2024arxiv), and TAP (TAP2024nips) often use other LLMs to generate or refine these templates.
-Concurrent work like StructTransform (structtransform2025arxiv) also targets structured generation, aligning with our findings.
+Methods like MasterKey (Deng et al., 2024), LLMFuzzer (Yu et al., 2024), AutoDAN (Liu et al., 2024a), PAIR (Chao et al., 2024b), and TAP (Mehrotra et al., 2024) often use other LLMs to generate or refine these templates.
+Concurrent work like StructTransform (Yoosuf et al., 2025) also targets structured generation, aligning with our findings.
 Linguistics-based Jailbreaks conceal malicious intent within seemingly benign inputs using linguistic properties.
-For example, DrAttack (drattack2024emnlpfinding) decomposes and reconstructs prompts, while Puzzler (puzzler2024aclfinding) uses combinations of diverse clues to bypass safety mechanisms.
-Encoding-based Jailbreaks obfuscate malicious prompts by translating them into less common languages (Zulu (zulu2024arxiv)) or encoding formats like Base64 (base64attack2023nips), exploiting the mismatch between the model’s high-level capabilities and its restricted safety alignment as identified by foundational work (wei2023jailbroken).
-Output-based Jailbreaks is an emerging category.
-White-box methods like EnDec (enforcedecode2024acl) and JailMine (jailmine2024arxiv) directly manipulate model logits or token probabilities to construct harmful outputs.
-Closely related to our work, APT (APT2025arxiv) uses the GuidedRegex feature in structured output APIs to build a prefix tree that explicitly bans refusal tokens, forcing a harmful response.
+For example, DrAttack (Li et al., 2024a) decomposes and reconstructs prompts, while Puzzler (Chang et al., 2024) uses combinations of diverse clues to bypass alignment.
+Encoding-based Jailbreaks obfuscate malicious prompts by translating them into less common languages like Zulu (Yong et al., 2024) or encoding formats like Base64 (Wei et al., 2023a), exploiting the mismatch between the model’s high-level capabilities and its restricted safety alignment as identified by foundational work (Wei et al., 2023b).
+Output-based Jailbreaks form an emerging category, which we formalize as Constrained Decoding Attacks (CDA).
+Early white-box methods like EnDec (Zhang et al., 2024) and JailMine (Li et al., 2024b) directly manipulated model logits to enforce harmful generation.
+In the black-box setting, APT (Li et al., 2025b) exploits GuidedRegex to iteratively block refusal tokens via a prefix tree.
+However, a major limitation of these approaches is their inefficiency, often requiring iterative interactions with the target model.
+In contrast, our proposed attacks are highly efficient, achieving jailbreaks with $O(1)$ query complexity (a single query) to the target LLM.
+While DictAttack leverages LLM assistance for dictionary construction, we emphasize that this process is offline, model-agnostic, and semantically benign: it generates harmless synonyms to construct a dictionary, without requiring any adversarial feedback or access to the target model’s internal states.
 7. Discussions
 Given the significant vulnerabilities revealed by CDAs, we discuss the implications for the LLM safety landscape and propose potential mitigations.
-As shown in Figure 6, current auditing practices focus on two phases: input and output auditing (outauditing2024arxiv).
+As shown in Figure 6, current auditing practices focus on two phases: input and output auditing (Bell and Fonseca, 2024).
 Input Auditing.
@@
 However, CDAs easily bypass this by hiding malicious intent within the control-plane grammar while keeping the data-plane prompt benign.
-While auditing the grammar alongside the prompt (Dual-Plane Guard) could mitigate this, it remains impractical for complex, interleaved multi-turn attacks due to high reasoning costs and the limited capability of lightweight guard models.
+While auditing the grammar alongside the prompt (Dual-Plane Guard) could mitigate this, it remains impractical for interleaved multi-turn attacks due to high auditing costs and the limited capability of lightweight guard models.
+Output Auditing.
+Output auditing filters generated content but faces steep challenges in latency, cost, and false positives (Mai et al., 2025).
+Consequently, major APIs like openai and gemini often omit it.
+Beyond practical limits, output auditing has a conceptual ceiling against CDAs: as we illustrate with BenignEnumAttack (Figure 9), an attack response (red flow) can be made structurally indistinguishable from a legitimate StrongREJECT-style safety audit (blue flow), forcing any purely post-hoc auditor into either a false positive on legitimate safety research or a false negative on the jailbreak.
 Figure 9.
-BenignEnumAttack (red) and a benign redteaming Q-A auditing process (blue).
-Existing output auditing methods cannot identify their differences, so that they will be refused or returned together, causing either false positive or false negative.
-Output Auditing.
-Output auditing filters generated content but faces steep challenges in latency, cost, and false positives (youcant2025www).
-Consequently, major APIs like openai and gemini often omit it.
-Beyond practical limits, we question if output auditing is a “silver bullet”.
-To test this, we developed BenignEnumAttack (Figure 9), which mimics a benign structured output (e.g., a StrongREJECT report) while embedding harmful content.
-This creates a semantic deadlock: since the jailbreak and a legitimate safety audit appear structurally identical, an auditor must either trigger a false positive (refusing safety research) or a false negative (leaking the jailbreak).
-This indicates that output-based defense has fundamental conceptual limitations beyond mere implementation costs.
-Mitigation Strategies. To secure the control plane, we propose several strategies across different layers:
+BenignEnumAttack (red) and a benign red-teaming Q-A auditing process (blue).
+Existing output auditing methods cannot identify their differences, causing either false positive or false negative.
+Mitigation Strategies. We sketch four directions for securing the control plane:
   1. (1)
-    Lightweight Heuristics: Filters could flag suspicious Literal-to-Logic Ratios or Semantic Mismatches between prompts and schemas.
-However, such heuristics face similar false positive risk in Figure 9.
-For example, a legitimate tool for generating a technical glossary or a translation table would naturally contain many string literals and might trigger these filters.Ensuring safety without breaking legitimate utility remains a complex trade-off that requires more robust and context-aware auditing.
+    Lightweight Heuristics: flag suspicious literal-to-logic ratios or prompt-schema semantic mismatches. Cheap, but inherently subject to the false-positive trade-off illustrated by Figure 9.
   2. (2)
-    Grammar-Plane Auditing: Direct auditing of JSON Schemas can mitigate basic attacks like EnumAttack. However, cross-plane coordination for DictAttack remains a high-cost reasoning challenge.
+    Grammar-Plane Auditing: directly audit user-supplied JSON Schemas. Sufficient for EnumAttack; cross-plane coordination for DictAttack remains a high-cost reasoning challenge.
   3. (3)
-    Context-Aware Token Attribution: Current output auditing processes cannot distinguish between user-prefilled content and model-generated content—a vulnerability motivated by BenignEnumAttack.
-A more robust approach would implement token-level provenance tracking during the generation process, enabling auditing systems to differentiate between tokens originating from enum constraints (or others specified by grammars) versus those freely generated by the model.
-This would enable more accurate detection of attempts to manipulate the generation trajectory through constrained fields, neutralizing threats like BenignEnumAttack.
+    Early-Stop Alignment (CB-style): train the model at the representation level to terminate or refuse harmful trajectories rather than relying on input-side filtering, as in Circuit Breakers (Zou et al., 2024).
+§§ 5.5 shows this materially suppresses single-vector attacks but does not close the gap once the prompt and grammar are composed (78% combined ASR), since the alignment side cannot reach the control plane.
   4. (4)
-    Safety-Preserving Constraints: We suggest whitelisting refusal tokens (e.g., “sorry”) to override user-defined grammars, ensuring internal safety alignment can trigger even within constrained outputs when jailbreak is absolutely detected.
-However, implementation challenges include balancing strict grammar adherence with safety requirements, as excessive safety tokens might disrupt valid structured outputs in certain contexts.
-  5. (5)
-    Integrated Safety Signaling: Co-design LLM with auditing signal tokens: we still whitelist some tokens in the constrained decoding process, but we whitelist special tokens that indicate potential auditing signals.
-For example, we can co-design LLMs to generate special auditing tokens (e.g., <unsafe>) when encountering harmful contexts, to enable real-time, constraint-agnostic monitoring.
+    Constraint-Aware Decoding: keep alignment-critical signals reachable under user grammars, e.g.
+unmaskable refusal tokens or constrained-decoding-transparent audit tokens (e.g., <unsafe>).
+Together with token-provenance tracking, this also neutralises BenignEnumAttack.
 8. Conclusion
@@
 Our evaluation across 13 models, including gpt-5 and gemini-2.5-pro, demonstrates that by weaponizing deterministic grammar constraints, attackers can bypass both internal alignment and external guardrails with near-perfect success.
-More importantly, our primary contribution, DictAttack, exposes a fundamental “semantic gap” in current defenses; by decoupling intent across space and time, it maintains a 75.8% ASR even against state-of-the-art jailbreak guardrails.
+More importantly, our primary contribution, DictAttack, exposes a fundamental “semantic gap” in current defenses: by realizing the control-to-semantic pipeline — the schema first commits malicious intent to the generation trajectory, after which the model fluently completes it — DictAttack maintains a 75.8% ASR even against SOTA jailbreak guardrails.
 By revealing this previously unexplored attack surface, our work contributes to developing more comprehensive security paradigms for LLMs that address safety at all stages of LLM generation.
 Ethical Considerations
-Our work identifies significant security vulnerabilities in constrained decoding mechanisms for LLMs by demonstrating how guided grammar manipulation can bypass safety guardrails.
-We empirically show that state-of-the-art models remain vulnerable to Constrained Decoding Attacks despite robust safety alignment efforts.
-The security implications are particularly concerning as CDAs exploit a fundamental gap in existing safety architectures, which primarily focus on input filtering rather than the generation process itself.
-By publishing this research, we aim to enable the security community to address these vulnerabilities before they can be widely exploited.
-Since CDAs target low-level decoding processes exposed through structured output features, mitigating these attacks presents unique challenges distinct from traditional jailbreak defenses.
-We have acknowledged our findings to affected model providers like OpenAI and Gemini and propose several mitigation strategies that could reduce risk without compromising legitimate structured output functionality.
-We believe our analysis provides crucial insights for developing more comprehensive safety mechanisms that protect the entire generation pipeline rather than just input and output boundaries.
+This paper exposes a previously underexplored attack surface — the LLM control plane — and shows that state-of-the-art models remain vulnerable to Constrained Decoding Attacks despite strong safety alignment.
+We disclosed our findings to OpenAI and Google (Gemini) in early 2025; the embargo has since passed.
+We additionally notified the maintainers of xgrammar, who acknowledged the vulnerability.
+Mitigations deployed inside closed-source GPT/Gemini stacks are not visible to us.
+We propose mitigation strategies that protect the full generation pipeline without disabling legitimate structured-output functionality, so that the community can close this gap before adversaries can exploit it.
+Open Science
+The project page here hosts self-contained proofs of concept and links to the full evaluation harness and run logs, which are available as a separate artifact under academic-only gated access; verified researchers may request access from the corresponding author.
+All datasets are open-access on Hugging Face.
+Because LLM evaluations involve sampling and model-based judging, results may vary slightly across runs.
+Acknowledgments
+This work is supported by the National Key Research and Development Program of China (24YFB4505603), the National Natural Science Foundation of China (62232015, 62302479, U23B2020), and the CAS Pioneer “Hundred Talents Program” (Category B, E545030000).
+We would like to thank anonymous reviewers for their insightful suggestions to improve this work.
+This paper was edited for grammar using Google Gemini (Gemini Team Google, 2023).